### v2

# Adaptive Parametric Activation

Abstract
The activation function plays a crucial role in model optimisation, yet the optimal choice remains unclear. For example, the Sigmoid activation is the de-facto activation in balanced classification tasks, however, in imbalanced classification, it proves inappropriate due to bias towards frequent classes. In this work, we delve deeper in this phenomenon by performing a comprehensive statistical analysis in the classification and intermediate layers of both balanced and imbalanced networks and we empirically show that aligning the activation function with the data distribution, enhances the performance in both balanced and imbalanced tasks. To this end, we propose the Adaptive Parametric Activation (APA) function, a novel and versatile activation function that unifies most common activation functions under a single formula. APA can be applied in both intermediate layers and attention layers, significantly outperforming the state-of-the-art on several imbalanced benchmarks such as ImageNet-LT, iNaturalist2018, Places-LT, CIFAR100-LT and LVIS and balanced benchmarks such as ImageNet1K, COCO and V3DET. The code is available at https://github.com/kostas1515/AGLU.
Keywords:
Activation function Long-tailed learning
1 Introduction
Image recognition has witnessed tremendous progress over the last years due to the use of deep learning, large image datasets such as ImageNet1K [17] and advances in model architectures [19, 28], learning algorithms [24, 69, 25], activation layers [31, 21, 27] and normalisation techniques [4, 40]. In this work, we focus on the activation layer of the network.
In balanced image classification works, it was empirically shown that if the activation function is close to the real data distribution then the model converges faster because the learning objective becomes easier [27, 31]. Based on this, the GELU [31] and the PRELU [27] were proposed as alternatives to the commonly used RELU [21] and they were utilised inside the model’s layers to activate the intermediate features of the network. Similarly, in imbalanced image classification, many works have empirically shown that the Sigmoid or the Softmax activation functions, are inappropriate and using another activation function increases the performance [71, 34, 63, 1]. Based on that, the Gumbel activation [1] and the Balanced Softmax [71] activation were proposed and they were used inside the classification layer to predict the classes. In contrast to balanced classification, these works focused only on the classification layer and they disregarded the importance of the intermediate activations. To this end, there is no principled way of choosing the right activation function, and usually, practitioners use the best activation function, according to the task, through cross-validation or parameter-tuning.
In this work, we focus on this problem. First, we theoretically show that the activation function enforces a prior belief of how the data is distributed and therefore it acts as an initialisation point. In practise, a good initialisation point enhances the convergence, therefore, having an appropriate activation function can increase the performance. Second, we study the impact of the activation function on balanced and imbalanced classification from two perspectives, i.e. the classification layer, and the intermediate layers.
Our findings show that the classification logit distribution of a pretrained model heavily depends on the degree of data imbalance. For example, in balanced training, the classification logits align better with the Logistic distribution, while in imbalanced learning they align better with the Gumbel distribution. Regarding the intermediate layers, we study the channel attention layer as an example and we find that it is also affected by the degree of data imbalance. We find that, in balanced learning, the channel attention is robust for all classes, however in imbalanced training the channel attention enhances more the frequent classes than the rare classes.
To this end, we empirically show that the commonly used Sigmoid activation function cannot generalise for both balanced and imbalanced learning, because it is non-parametric and does not align with the imbalanced data distribution. Motivated by this, we develop a novel Adaptive Parametric Activation (APA) function. APA allows the model to align its activations to both the balanced and imbalanced data distributions and reach great performance in both tasks. APA has several benefits, it unifies most previous activations functions such as the Sigmoid, the Gumbel [1], the RELU [21], the SiLU [31] and the GELU[31] under a common formula. Also, it uses two learnable parameters that allow the network to select the best activation function during optimisation enlarging the model’s capacity. Moreover, our APA is versatile, it can be used as a direct replacement to RELU, or it can replace the Sigmoid activation function inside the attention mechanism boosting the performance significantly and consistently. Finally, APA can be generalised to both imbalanced and balanced classification and detection tasks and surpass the state-of-the-art (SOTA). Our contributions are:
  • We demonstrate the importance of the activation function in balanced and imbalanced data distributions, through statistical analysis;
  • We propose the novel APA function that unifies most common activation functions under a single formula;
  • We have validated the efficacy of APA on a range of long-tailed benchmarks, including ImageNet-LT [58], iNaturalist18 [84], Places-LT [58], CIFAR100-LT [8], LVIS [22] and balanced benchmarks such as ImageNet1k [17], COCO [55] and V3Det [88] largely surpassing the state of the art.
2 Preliminaries
Activation function. First, we show the importance of the activation function in model optimisation, following [1]. Let’s consider the example of binary classification, where $z$ is the input, $y\in\{0,1\}$ is the target class and $f(z)=W^{T}z+b$ is the classification network. The target class and the input $z$ are related as:
$y=\begin{cases}1,\;\;\text{if}\;\;f(z)+\epsilon>0\\ 0,\;\;\text{otherwise}\end{cases}$ (1)
where $\epsilon$ is the error, that is a random variable and it is distributed according to the real data distribution. In this example, the classification boundary is set to 0, however, it can be adjusted by the network’s bias $b$ during optimisation. The probability of the class $P(y=1)$ is:
$P(y=1)=P(f(z)+\epsilon>0)=P(\epsilon>-f(z))=1-F(-f(z))$ (2)
where $F$ is the cumulative distribution function. During network optimisation, if one chooses the Sigmoid activation function $\sigma(z)=(1+\exp(-z))^{-1}$, then the prediction $\bar{y}$ is obtained using $\bar{y}=\sigma(f(z))$ and the probability $P(\bar{y}=1)$ is:
$P(\bar{y}=1)=\frac{1}{1+\exp(-f(z))}=F_{\text{logistic}}(f(z);0,1)=1-F_{\text{logistic}}(-(f(z);0,1)$ (3)
Comparing Eq. 2 and Eq. 3, it is shown that when we use the Sigmoid, we assume that the error term $\epsilon$ follows the standard Logistic distribution. If we use Gumbel activation then we assume that the error follows the Gumbel distribution [1] and, in general, any activation function assumes a different distribution of $\epsilon$.
For this reason, the activation function can be seen as an initialisation point, or a prior belief of how the real data are distributed. If the prior is good, then the learning objective becomes easier and the performance is increased as it was shown empirically by many past studies [27, 1, 21, 77].
2.1 Balanced versus Imbalanced learning
In this subsection, we perform a statistical analysis on the empirical distributions of the logits and the intermediate activations, to understand the importance of the activation function in balanced and imbalanced learning.
Classification logits. We train a MaskRCNN [26] ResNet50[28] on LVIS dataset [22], which is a highly imbalanced object detection dataset containing 1203 classes. After the model has converged, we perform inference and store the predicted classification logits, i.e., $f(z_{y})$ for every class $y$. Next, we perform histogram binning on the logits and we visualise the empirical distributions for the rare class puffin and the frequent class glove in Fig. 1 (b) and (c) respectively.
The rare class has a negative average logit value, because it is dominated by the frequent classes, which have higher average logit values as shown in (c). Regarding its distribution, it resembles the Gumbel distribution, because it has a heavy right tail and it is skewed. To quantify that, we calculate the statistical distance, using the Kolmogorov-Smirnov (KS) test [62], between all empirical class distributions and the Gumbel and Logistic theoretical distributions, shown in (a) and (e). As shown in (d), most classes have smaller distance to the Gumbel distribution. We repeat this test, using ResNet50 trained on the balanced ImageNet1K. As shown in (f) and (g), the classification logits are different this time, for example most average logit values are centered around zero and they are less skewed. In general, the logit distributions are closer to the logistic distribution as shown in (h). This explains why the Sigmoid activation achieves better performance in the balanced classification task and worse performance in imbalanced classification [1, 71, 34, 63]. Next, we show that data imbalance also affects the intermediate layers, by studying the channel attention as an example.
Figure 1: Top: In imbalanced learning, the logit distributions are more skewed and they have a smaller KS distance to the Gumbel than the Logistic distribution as shown in (d). Bottom: In balanced learning, the logit distributions are less skewed and they align better with the Logistic, than the Gumbel distribution, as shown in (h).
Attention layer. Attention mechanism for input $X\in\mathbb{R}^{H\times W\times C}$ re-weights the input features $X$ by applying an attention function $A(X)$, i.e., $X^{\prime}=A(X)\otimes X$. For example, in Channel Attention (CA) [37], $X^{\prime}=\sigma(\text{MLP}(\text{GAP}_{c}(X)))\otimes X$, where $\text{GAP}_{c}\in\mathbb{R}^{1\times 1\times C}$ is Global Average Pooling and $\otimes$ is the element-wise product. In this case, the attention function is $A_{CA}(X)=\sigma(\text{MLP}(\text{GAP}_{c}(X)))$.
Figure 2: Visualisations of channel attention (A). In (a) the attention signals when training with imbalanced ImageNet and balanced ImageNet have similar variance in the first layer but completely different in the last most semantic layer in (b). The variance with ImageNet-LT training drops to zero for deeper layers as shown in (c), because the attention promotes only few frequent classes. In (d) and (e) the entropy of channel attention is smaller for the rare classes than the frequent classes in both i-Naturalist18 and ImageNet-LT training. In (f) and (g) the channel responses are smaller for the rare classes than the frequent ones for i-Naturalist18 and ImageNet-LT respectively.
Balanced vs Imbalanced channel attention. We train SE-ResNet50 [37] models (SE-R5) on balanced ImageNet-1K and imbalanced ImageNet-LT. After training the models, we analyse the average channel attention signals for a random batch of 128 test images. Fig. 2-a shows the output $A_{CA}$, in the first layer of SE-R50, Fig. 2-b shows the output, in the last layer, and Fig. 2-c shows the variance of $A_{CA}$ across all layers. As shown in Fig. 2-a, the attention is similar in the first layer and it is different in the last layer, as shown in Fig. 2-b. As displayed in Fig. 2-c, the attention variance with ImageNet1K (blue-curve) is larger and the attention reweights all channels and affects all classes. In contrast, for the imbalanced case (orange-curve), the attention signal has small variance, indicating that it is biased to some classes.
Layer-wise analysis. This phenomenon is most prevalent in the last attention layer, which is the most semantic. To quantify which attention layers are affected the most, we use entropy as a measure of signal complexity. Since the channel attention produces a probabilistic weighting vector via the Sigmoid activation, we calculate the total entropy of channel attention, for a layer $l$, as the sum of the binary channel distribution entropies as follows:
$\begin{split}E_{l}=-\frac{1}{C}\sum_{i=1}^{C}[(A_{CA}(X_{l})\log(A_{CA}(X_{l}))+(1-A_{CA}(X_{l})\log(1-A_{CA}(X_{l}))]\end{split}$ (4)
When the layer’s entropy is closer to zero, the channel attention signals are closer to 1 and they do not affect the original features for that layer, i.e., $X^{\prime}=1\otimes X$. If the layer’s attention entropy is closer to one, then the channel attention signals are informative, as they affect the signal i.e., $X^{\prime}=A(X)\otimes X$.
To investigate the complexity of the attention signal, we propagate two batches of 64 test images that contain only frequent and only rare classes respectively, through the pretrained SE-R50 and measure the attention entropy of $A_{CA}$. In Fig. 2 (d) and (e), the average entropy is similar for both frequent and rare classes for all layers except for the last layer which is the most semantic. The blue curve, that corresponds to rare class channel attention, has lower entropy than the orange curve that corresponds to frequent class channel attention in the last layer for both i-Naturalist-18 and ImageNet-LT. This indicates that channel attention produces simpler signals for the rare classes and more complex signals for the frequent classes. Finally, in (f) and (g), we show that the average channel responses are smaller when the inputs are rare classes, than frequent classes, which explains why the network cannot model the rare classes effectively. In conclusion, this analysis shows that data imbalance affects the quality of the activations inside the intermediate layers and it highlights the limitation of the Sigmoid activation to model the rare classes.
Figure 3: Our APA unifies most activation functions under the same formula.
3 Method
As shown in the previous section, the degree of data imbalance affects both the classification logits and the intermediate layers. While it is possible to perform a statistical analysis and select the appropriate activation for the classification layer, this is difficult to do for all layers of the network, because first, the data distributions inside the layers dynamically change during training  [4, 40], and secondly, there is no one-to-one correspondence between classes and intermediate channels, which hinders attribution. For this reason, we propose the Adaptive Parametric Activation (APA):
$\begin{split}\eta_{ad}(z,\kappa,\lambda)=(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}\\ \end{split}.$ (5)
APA can adjust its activation rate, dynamically, according to the input’s distribution using two parameters $\kappa$ and $\lambda$, that can be learned during optimisation. $\kappa\in\mathbb{R}$ is the gain parameter that controls the function’s sensitivity. $\lambda\in(0,\infty)$ is the asymmetrical parameter that controls the function’s response rate to positive and negative inputs, allowing the model to have different learning degrees when the input is positive or negative. This function is also known as Richard’s curve [73] and it unifies the most common activation functions.
For example, if $\kappa=\lambda=1$ then APA becomes the Sigmoid activation that has a symmetric response rate for both positive and negative inputs and it is successful for balanced classification tasks. If $\kappa=1,\lambda\xrightarrow[]{}0$ then APA becomes the Gumbel activation that has an asymmetric response rate and it is successful for long-tailed instance segmentation tasks [1]. This behaviour is shown in Fig. 3 left and middle. Based on Eq. 5, we also define the Adaptive Generalised Linear Unit (AGLU):
$\begin{split}AGLU(z,\kappa,\lambda)=z\cdot\eta_{ad}(z,\kappa,\lambda)\\ \end{split}$ (6)
Figure 4: AGLU derivatives with respect to $\kappa$ (top), $\lambda$ (middle) and $z$ (bottom).
AGLU has many interesting properties, for example, if $\kappa=\lambda=1$ then AGLU becomes the Sigmoid Linear Unit (SiLU) [31]. If $\kappa=1.702,\lambda=1$, then it becomes the Gaussian Error Linear Unit (GELU) [31]. If $\kappa\xrightarrow[]{}\infty$, then AGLU becomes ReLU[21] and if $\lambda\xrightarrow[]{}\infty$, then AGLU becomes the identity function, as shown in Fig. 3-right.
In other words, the $\kappa$ parameter controls the RELU-ness and the $\lambda$ parameter controls the leakage. Also, AGLU could be seen as a smoother version of PRELU, because $AGLU(z,1,\lambda\xrightarrow[]{}\infty)=PRELU(z,1)$. We compare $AGLU$ to most common activation functions, in more detail, in Table 1. The derivative of AGLU with respect to $\kappa$ is:
$\frac{\partial AGLU(z,\kappa,\lambda)}{\partial\kappa}=z^{2}\cdot\frac{\eta_{ad}(z,\kappa,\lambda)}{\lambda+\exp(\kappa z)}$ (7)
the derivative of AGLU with respect to $\lambda$ is:
$\frac{\partial AGLU(z,\kappa,\lambda)}{\partial\lambda}=-\frac{z}{\lambda}\cdot\frac{\eta_{ad}(z,\kappa,\lambda)}{\lambda+\exp(\kappa z)}$ (8)
and the derivative of AGLU with respect to $z$ is:
$\frac{\partial AGLU(z,\kappa,\lambda)}{\partial z}=\kappa z\cdot\frac{\eta_{ad}(z,\kappa,\lambda)}{\lambda+\exp(\kappa z)}+\eta_{ad}(z,\kappa,\lambda)$ (9)
The proofs of the derivatives are shown in the Appendix. The derivatives of AGLU are shown in Fig. 4. Using various $\kappa$ and $\lambda$ combinations AGLU has drastically different behaviour and this enhances the capacity of the network and achieves good performance as shown in the experiments.
Table 1: Comparison of different activation functions.
Name Formula Range
RELU [21] $\eta(z)=\max(0,z)$ $(0,\infty)$
Gaussian Unit [31] $\eta(z)=z\sigma(1.702z)$ $(-0.17,\infty)$
Sigmoid Unit [31] $\eta(z)=z\sigma(z)$ $(-0.28,\infty)$
Mish [64] $\eta(z)=z\tanh(\ln(1+\exp(z))$ $(-0.31,\infty)$
PRELU [27] $\eta(z,\kappa)=\max(0,z)+\kappa\min(0,z)$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
ELU [11] $\eta(z,\kappa)=\max(0,z)+\kappa(\exp(\min(0,z))-1)$ $(-\kappa,\infty)$
AGLU (ours) $\eta(z,\kappa,\lambda)=z\cdot(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
In the end, our APA is versatile because it can be used not only as an error unit, replacing RELU, but also as an activation function inside the attention mechanism, replacing the Sigmoid activation.
4 Experiment Setup
Datasets. We use CIFAR100-LT [8] with exponential imbalance factor of 100 and 10, ImageNet-LT [58], Places-LT [58] and iNaturalist2018 [84] following the common long-tailed classification protocol. We report our results using top-1 accuracy on the balanced test sets, to fairly evaluate all classes. For ImageNet-LT, we split the categories according to their class frequency in the training set, into Many ($>$100 images), Medium (20$\sim$100 images) and Low ($<$20 images) and do per-group evaluation following [58]. Also, we use the LVISv1 [22] instance segmentation dataset, which has 100K training images and 1203 classes, that are grouped according to Frequent ($>$100 images), Common (10$\sim$100 images) and Rare ($<$10 images) classes. For this dataset, we report mask average precision $AP$, bounding box average precision $AP^{b}$ and $AP^{r}$, $AP^{c}$ and $AP^{f}$ which is mask average precision for rare, common and frequent classes respectively. Regarding balanced training, we perform experiments on ImageNet1K [17], COCO [55] and the recently proposed V3Det[88], which is a challenging large scale detection dataset with 13K classes and 243K images. We report top-1 accuracy for ImageNet1K, $AP^{b}$ and $AP^{m}$ for COCO and $AP^{b}$ for V3Det.
Table 2: Top-1 accuracy (%) on ImageNet-LT test set. E denotes ensemble.
Method Backbone Many Medium Few Average
MiSLAS [111] R50 [28] 61.7 51.3 35.8 52.7
KCL [43] 61.8 49.4 30.9 51.5
TSC [53] 63.5 49.7 30.4 52.4
RIDE (3E)+CMO[68] 66.4 53.9 35.6 56.2
DOC [86] 65.1 52.8 34.2 55.0
CC-SAM [116] 61.4 49.5 37.1 52.4
Our Baseline SE-R50 [37] 66.2 53.1 37.1 56.0
APA* (ours) 67.5 54.3 39.3 57.4
APA* + AGLU (ours) 68.3${}^{\textbf{+1.9}}$ 54.8${}^{\textbf{+0.9}}$ 39.4${}^{\textbf{+2.1}}$ 57.9${}^{\textbf{+1.7}}$
RIDE (4E) [95] X50 [100] 68.2 53.8 36.0 56.8
SSD [52] 66.8 53.1 35.4 56.0
BCL [117] 67.9 54.2 36.6 57.1
CNT  [67] 63.2 52.1 36.9 54.2
ALA  [110] 64.1 49.9 34.7 53.3
ResLT  [13] 63.6 55.7 38.9 56.1
ABC-Norm [36] 60.7 49.7 33.1 51.7
RIDE (3E)+CMO+CR [60] 67.3 54.6 38.4 57.4
LWS+ImbSAM [115] 63.2 53.7 38.3 55.3
Our Baseline SE-X50 [37] 67.9 53.0 37.7 56.7
APA* (ours) 68.9 55.4 39.4 58.4
APA* + AGLU (ours) 69.8${}^{\textbf{+1.6}}$ 55.70.${}^{\text{0.0}}$ 41.1+2.${}^{\textbf{+2.2}}$ 59.1${}^{\textbf{+1.7}}$
Implementation Details We primarily use Squeeze and Excite [37] as our baseline with ResNet-32 [28] for CIFAR-LT, ResNet50 for iNaturalist, ResNet50 and ResNext50 [100] for ImageNet-LT and ResNet152 for Places-LT, which has been pretrained on ImageNet1K according to [58]. For LVIS, we use SE-Resnets with MaskRCNN [26], FPN [54], Normalised Mask [89], RFS sampler [22] and GOL loss [1] as a baseline. For V3Det, we use SE-ResNet50 with FasterRCNN [72], FPN and Wrapped Cauchy classifier [23]. All baselines use bag of tricks [109], and strong training techniques that we discuss and ablate in the Appendix. Our motivation for using bag of tricks is two-fold, first, it pushes the performance even further and secondly, it showcases the generalisability of our work.
For our attention models, we replace the Sigmoid with APA, and we further use LayerNorm [4] and attention dropout [32] with $p=0.1$, for all datasets except iNaturalist. We denote this configuration as APA* in our Tables and we ablate its components. For our AGLU models, we simply use it in-place of RELU.
5 Results
Table 3: (a) Results of APA* and AGLU for iNaturalist and Places-LT. (b) Results of APA* and AGLU for CIFAR100-LT with imbalance 10 and 100.
Method iNat18 PlacesLT
DisAlign [108] 70.6 39.3
MisLAS [111] 71.6 40.4
LADE [34] 70.0 38.8
ALA [110] 70.7 40.1
TSC [53] 69.7 -
CNT [67] - 39.2
WD+MaxNorm [3] 70.2 -
DOC [86] 71.0 -
BCL[117] 71.8 -
ResLT [13] 70.5 39.8
IIF [2] - 40.2
ABC Norm [36] 71.4 -
LWS+ImbSAM[115] 71.1 -
CC-SAM[116] 70.9 40.6
AREA [10] 68.4 -
Our Baseline with SE 71.3 40.5
APA* (ours) 72.3 41.3
APA* +AGLU (ours) 74.8${}^{\textbf{+3.0}}$ 42.0${}^{\textbf{+1.4}}$
Method 10 100
BALMS [71] 63.0 50.8
RIDE (4E) [95] - 49.4
ACE (4E) [7] - 49.6
DiVE [29] 62.0 45.4
SSD [52] 62.3 46.0
MisLas [111] 63.2 47.0
HSC [91] 63.1 46.7
LADE [34] 61.7 45.4
ResLT (3E) [13] 63.7 49.7
TLC (4E) [50] - 49.8
TSC [53] 59.0 43.8
RIDE+CMO [68] 60.2 50.0
CC-SAM [116] - 50.8
RIDE+CMO+CR[60] 61.4 50.7
AREA [10] 60.8 48.9
Our Baseline with SE 65.2 50.9
APA* (ours) 65.7 51.9
APA*+AGLU(ours) 66.8${}^{\textbf{+1.6}}$ 52.3${}^{\textbf{+1.4}}$
Long-tailed Classification Benchmark. We compare APA* against ensemble and fusion models [95, 13, 50, 7], margin adjustment [34, 108, 110], contrastive learning [91, 53, 43], knowledge transfer [67], knowledge distillation [29, 52], decoupled methods [111, 2, 108, 36], sharpness aware minimisation [115, 116, 60] and data augmentation [111, 68].
On ImageNet-LT, as shown in Table 2, our baseline models with bag of tricks reach the state-of-the-art (SOTA) for both SE-ResNet50 (R50) and SE-ResNeXt50 (X50). We want to point out that most performance comes from the bag of tricks and not the SE module, as shown, in detail in the Appendix.
Our APA* outperforms the SE-R50 baseline by $1.4$ percentage points (pp) on average, by $1.3$pp on frequent categories, $1.2$pp on medium and $2.2$pp on few classes. Most importantly, it increases the performance of both frequent and rare classes, which is a unique advantage compared to the previous works. Additionally, our APA* exceeds RIDE with 3 Experts (3E) and CMO [68] by $1.2$pp on average and by $3.7$pp on the rare classes using a single model. When AGLU is combined, it pushes the performance of APA* R50, by $0.5$pp on average, by $0.8$pp on the frequent, $0.5$pp on the medium and $0.1$pp on the few classes.
APA* with X50 increases the SE-X50 baseline by $1.7$pp on average, $1.0$pp the many classes, $1.4$pp the medium classes and $1.7$pp the few classes. Furthermore, it outperforms RIDE(3E)+CMO+CR [60] by $1.0$pp on average, $1.6$pp on frequent classes, $0.8$ on common and $1.0$pp on rare classes using a single model. When AGLU is combined to APA* X50, it adds $0.7$pp on average, $0.9$pp on the frequent categories, $0.3$pp on the medium, and $1.7$pp on the few classes.
Table 4: For all ablations we use ImageNet-LT. In (a) we compare adaptive activation to other learnable activation functions using SE-ResNet50. In (b) we show that adaptive activation and AGLU generalise to other attention mechanisms. In (c) we show that APA* can be combined with classifier learning techniques. In (d) we show that APA* works effectively with deeper ResNets. In (e) we show the components of APA*. In (f) we compare AGLU with other activation functions.
Activation Many Med. Few Avg
Sigmoid 66.2 53.1 37.1 56.0
with Temp 65.9 53.8 40.3 56.6
Gumbel 66.2 53.2 39.3 56.3
with Temp 66.9 53.4 39.7 56.7
[…]
Attention type Avg
Spatial [98] 54.8
+APA 55.2
+APA + AGLU 56.4
Spatial + Channel  [98] 55.6
+APA 56.9
+APA + AGLU 57.1
Method SE APA*
CE 51.7 52.9
PC-Softmax [34] 56.0 57.4
cRT [44] 55.6 56.4
Decoupled-DRW  [8] 55.1 56.6
BSCE [71] 56.0 57.0
(a) APA comparison.
(b) Attention types w/ R50.
(c) Classifier learning.
Method Many Med. Few Avg
SE-R101 67.5 53.3 37.6 56.7
[…]
SE-R152 68.0 54.4 39.8 57.6
[…]
SE APA Dropout LayerNorm Avg
55.0
56.0
57.0
57.3
57.4
Activations Avg
ReLU 57.4
PReLU [27] 54.8
ELU [11] 52.6
Mish [64] 57.4
GELU [31] 57.5
SiLU [31] 57.1
AGLU 57.9
(d)Deeper Networks.
(e) APA* ablation.
(f) AGLU comparison.
On iNaturalist18, Places-LT and CIFAR100-LT, our SE baselines with bag of tricks again reach the SOTA, as shown in Table 3-a and b.
Regarding, iNaturalist18, APA* improves the SE-baseline by $1.0$pp. It also outperforms BCL [117] by $0.5$pp, ResLT by $1.8$pp and LWS+ImbSAM [115] by $1.2$pp. AGLU further enhances the performance by a staggering $3.5$pp compared to SE-baseline, which is a significant increase.
On Places-LT, our APA* also increases the performance of the SE baseline by $0.8$pp. Moreover, it surpasses MisLAS [111] by $0.9$pp and CC-SAM [116] by $0.7$pp. When AGLU is combined with APA*, it further increases the performance by $0.7$pp reaching $42.0\%$.
As shown in Table 3-b, on CIFAR100-LT, APA* improves the performance by $0.5$pp and $1.0$pp compared to SE baseline using an imbalance factor of 10 and 100 respectively. Furthermore, compared to RIDE+CMO [68], CC-SAM [116] and RIDE+CMO+CR [60], APA* achieves $1.9$pp, $1.1$pp and $1.2$pp higher accuracy respectively, under an imbalance factor of 100 using a single model. Finally AGLU, further boosts the performance of APA* by $1.1$pp and $0.4$pp for imbalance factor of 10 and 100 respectively.
Activation Ablation Study. We compare APA, without the Dropout and Layernorm, against different activation functions such as the Sigmoid and the Gumbel [1] using the SE-ResNet50 baseline. Also, we implement Sigmoid and Gumbel variants with learnable temperature, to further understand their difference to our adaptive activation. As shown in Table 4-a, Sigmoid with temperature achieves slightly better performance for the rare classes, however it over-fits the frequent categories. In contrast, our APA achieves the best performance, increasing the overall performance by $0.3$pp, the frequent classes by $0.2$pp and the common categories by $0.4$pp compared to the second best Gumbel with temperature.
Generalisation to other Attention mechanisms. In Table 4-b, we show that APA can be combined with other attention mechanisms. Specifically, APA improves the performance of Spatial Attention by $0.4$pp and the Spatial-Channel Attention by $1.3$pp. AGLU further increases the performance of Spatial Attention by $1.2$pp and Spatial-Channel Attention by $0.2$pp.
Table 5: Comparisons on LVISv1.0 using MaskRCNN-FPN and 2x schedule.
Method Backbone $AP^{m}$ $AP^{r}$ $AP^{c}$ $AP^{f}$ $AP^{b}$
RFS  [22] R50 23.7 13.3 23.0 29.0 24.7
IIF [2] 26.3 18.6 25.2 30.8 25.8
Seesaw [89] 26.4 19.6 26.1 29.8 27.4
LOCE [20] 26.6 18.5 26.2 30.7 27.4
PCB+Seesaw [30] 27.2 19.0 27.1 30.9 28.1
ECM  [39] 27.4 19.7 27.0 31.1 27.9
GOL  [1] 27.7 21.4 27.7 30.4 27.5
ECM+GAP [107] 26.9 20.1 26.8 30.0 27.2
GOL (baseline) SE-R50 28.2 20.6 28.9 30.8 28.1
GOL+AGLU(ours) APA*-R50 29.1${}^{\textbf{+0.9}}$ 21.6${}^{\textbf{+0.2}}$ 29.6${}^{\textbf{+0.7}}$ 31.7${}^{\textbf{+0.6}}$ 29.0${}^{\textbf{+0.9}}$
RFS [22] R101 [28] 27.0 16.8 26.5 32.0 27.3
NorCal [65] 27.3 20.8 26.5 31.0 28.1
Seesaw [89] 28.1 20.0 28.0 31.8 28.9
GOL [1] 29.0 22.8 29.0 31.7 29.2
ECM [39] 28.7 21.9 27.9 32.3 29.4
PCB + Seesaw[30] 28.8 22.6 28.3 32.0 29.9
ROG [107] 28.8 21.1 29.1 31.8 28.8
GOL (baseline) SE-R101 29.7 23.0 29.9 32.5 30.0
GOL+AGLU(ours) APA*-R101 30.7${}^{\textbf{+1.0}}$ 23.6${}^{\textbf{+0.6}}$ 31.3${}^{\textbf{+1.4}}$ 33.1${}^{\textbf{+0.7}}$ 31.11.${}^{\textbf{1.1}}$
Combining APA* with Classifier Learning. APA* is an efficient module that can be easily combined with common classifier learning techniques such as margin adjustment [71], reweighting [8] and resampling [44]. As shown in Table 4-c, APA* consistently boosts the performance of all these methods.
Larger ResNets. We further train deeper models like ResNet-101 and ResNet-152 on ImageNet-LT and compare the SE and APA* attention methods. As Table 4-d shows, the APA* module enlarges the performance of all models consistently for both frequent, common and rare classes. Especially, APA* with ResNet101 promotes overall accuracy by $2.1$pp, boosting the rare categories by a significant $4.5$pp compared to SE-Resnet101.
Ablation of APA* components. APA* uses Channel attention, APA, Dropout and Layer-Norm and we show their effects in Table 4-e. The plain R50, without channel attention, achieves $55.0\%$ and SE increases its performance by $1.0$pp. APA increases the performance of the SE baseline by another $1.0$pp. Dropout increases the performance by $0.3$pp and LayerNorm enhances the performance by an additional $0.1$pp. In the Appendix, we show the full component ablation.
Comparison of AGLU. In Table 4-f, we compare AGLU to other commonly used activations functions, using APA* ResNet50 as backbone and switching the activation function of the intermediate layers. Our AGLU outperforms all the other methods and it is the best choice.
Long-Tailed Instance Segmentation. As the results suggest in Table 5, the SE-R50 backbone increases the overall mask performance compared to plain GOL-R50 by $0.5$pp but most improvement comes from the common and frequent categories while the rare categories are significantly reduced by $0.8$pp. In contrast, our APA* with AGLU-R50 improves the performance by $0.9$pp on average mask and bounding box precision, $1.0$pp on $AP_{r}$, $0.7$pp on $AP_{c}$ and $0.9$pp on $AP_{f}$ compared to SE-R50-GOL. Compared to GOL with SE-R101, APA* with AGLU also improves the performance by $1.0$pp on $AP$, $0.6$pp on $AP^{r}$, $1.4$pp on $AP^{c}$, $0.7$pp on $AP^{f}$ and $1.1$pp on $AP^{b}$. This highlights that our APA* and AGLU modules are robust for the rare classes and they outperform the previous SOTA in long-tailed instance segmentation.
Figure 5: a) APA* (orange curve) increases the attention variance by $0.04$ in the most semantic layer compared to the baseline (blue curve) and removes frequent category attention bias. b) APA* increases the attention entropy in the most semantic layer and retrieves rare class descriptors more efficiently than the baseline. c) Compared to the baseline (top), APA* produces larger entropy attention signals and makes correct rare class predictions. More visualisations are shown in Appendix.
Qualitative Results. We use APA* and Imagenet-LT for our qualitative analysis. In Fig. 5-(a) we show that APA* increases the variance of channel attention for the most semantic layer by $0.04$ compared to the baseline, making it diverse and informative for all channels. In Fig. 5-(b), we show that APA* increases the entropy of the attention signal for the deeper layers by $0.4\sim 0.6$ compared to the baseline, showing that our module produces informative signals that effectively attend to the rare classes. In Fig.-c 5-top, we show that the baseline channel attention of the rare classes, produces all-pass filters in other words, attention signals that have small entropy, i.e., $E=0.018$ for the bee house and $E=0.098$ for the warthog respectively. This hinders rare class learning and it results in misclassification. In contrast, our method in Fig. 5-bottom produces informative channel attention signals that have larger entropy i.e. $E=0.46$ and $E=0.91$ for the bee house and warthog classes respectively, allowing the model to retrieve rare class features and to make correct predictions. Visualisations of the learned $\kappa$ and $\lambda$ parameters are provided in the Appendix.
Generalisation to Balanced tasks. In Table 6-a, we show that APA*+AGLU increases the performance of SE-MaskRCNN-R50 by $0.7$pp on $AP^{b}$ and $AP^{m}$ on COCO. In Table 6-b, we show that APA*+AGLU increase the performance of SE-FasterRCNN-R50 by $2.9$pp and the performance of SE-Cascade-RCNN-R50 by $2.1$pp on the challenging V3Det, which are significant increases considering that this dataset has 13K classes. In Table 6-c, we show that our methods increase the performance for R50, SE-R50 and CBAM-R50, by $0.6$pp, $1.2$pp and $0.6$pp respectively. Finally in Table  6-d, we show that our modules increase the performance by $1.2$pp on SE-R50, $0.9$pp on SE-R101 and $0.5$pp on SE-R101, showing that they can generalise in balanced training.
Table 6: Resuls on balanced tasks. In (a) and (b), we combine APA*+AGLU on COCO dataset and V3Det dataset respectively. In (c) we combine AGLU and APA* to ResNet50, SE-ResNet50 and CBAM-ResNet50 in ImageNet1K. In (d) we show that APA*+AGLU generalises to deeper backbones in ImageNet1K.
Method $AP^{b}$ $AP^{m}$
MaskRCNN-R50 39.2 35.4
w/ SE 40.5 36.9
w/ APA*+AGLU 41.2 37.6
Method $AP^{b}$
FasterRCNN-R50 25.4
w/ SE 27.0
w/ APA*+AGLU 29.9
CascadeRCNN-R50 31.6
w/ SE 33.3
APA*+AGLU-CRCNN 35.4
Method top-1
ResNet50 [37] 76.9
w/ AGLU 77.5
SE-ResNet50 [37] 77.5
w/ APA*+AGLU 78.7
CBAM-ResNet50 [37] 78.3
w/APA*+AGLU 78.9
Method top-1
SE-R50 77.5
+APA* 77.9
+APA* + AGLU 78.7
SE-R101 79.4
+APA* 79.2
+APA* + AGLU 80.3
SE-R152 80.3
+APA* 80.5
+APA* + AGLU 80.8
(a)COCO
(b)V3Det
(c)ImageNet1k
(d) ImageNet1k
6 Related Work
Long-tailed image recognition. Long-tailed image recognition can be grouped according to representation learning and classifier learning techniques. Representation learning techniques improve the feature quality through rare class features generation [85, 87, 104], contrastive objectives [91, 15, 117, 74, 53, 43], ensemble or fusion models [95, 51, 14, 113, 50, 13, 7], knowledge distillation [51, 41, 29, 52], knowledge transfer [58, 67, 118] and data augmentation [111, 68, 101]. Classifier learning techniques enhance rare class classification through decoupled training [44, 108, 92, 47, 36], margin adjustment [63, 71, 34, 8, 39, 65, 89, 93, 110, 2], cost-sensitive learning [16, 46, 96, 38], resampling [9, 61, 76, 120, 22, 68, 33, 104], dynamic loss adaptation based on batch statistics [35, 80, 94], gradient statistics [79, 49], weight norms [93] and classification scores [20, 30]. These techniques are efficient, however they are applied either during the input phase or during the loss disregarding the intermediate channel activations.
Attention Networks. Attention networks, such as spatial attention [42], channel attention [37] and spatial-channel attention [98], are widely used in image recognition. Self-attention is the core mechanism of the Visual Transformer (ViT) architecture [19], and recently many improvements have been made to the Transformer architecture [81, 83, 56] and its training procedure [78, 6, 82].
In our work, we primarily used APA with channel attention models, but it also works with plain ResNets, Spatial and Spatial-Channel attention models. In the Appendix, we show results with self-attention models.
Activation functions The RELU function [21] dominates the landscape of deep image classification, and it is especially used inside the convolutional networks, while the GELU and SiLU [31] are more commonly used inside the transformer network [19] or the ConNext models [57, 97]. The PRELU function [27] is a generalisation of the RELU, because it linearly activates both positive and negative inputs and the ELU [11] is a follow-up work that enforces saturation on the negative inputs after some threshold. Our AGLU activation, generalises the RELU, the GELU and SiLU and it can also be seen as a smoother version of PRELU.
7 Conclusion
Our work highlights the impact of the activation function inside the model’s activations for balanced and imbalanced data distributions. We have empirically showed that the degree of data imbalance affects the logit distributions and the intermediate signals and we have showed that the commonly used Sigmoid activation function is unable to model the intermediate features. To this end, we have proposed an novel adaptive parametric activation that unifies most common activation functions under the same formula and we have tested it in several long-tail and balanced classification and detection benchmarks showing great generalisation.
Acknowledgements. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) project “ViTac: Visual-Tactile Synergy for Handling Flexible Materials” (EP/T033517/2).
Appendix 0.A Representations’ Quality
We evaluate the quality of the representations learned by the SE and APA* models using the recently proposed Neural Collapse framework [66].
Let $f_{k,j}\in\mathrm{R^{d}}$ be the features of the penultimate layer, $k=\{1,2,...,K\}$ the class, $n_{k}$ the number of samples in the class $k$ and $n=\sum_{k=1}^{K}n_{k}$ the total number of samples in the dataset. Then the global feature $f_{G}$ and class prototype $\bar{f_{k}}$ are:
$f_{G}=\frac{1}{n}\sum_{k=1}^{K}\sum_{j=1}^{n_{k}}f_{k,j}\;,\bar{f_{k}}=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}f_{k,j}$ (10)
The within-class covariance matrix $\Sigma_{W}\in\mathrm{R^{d\times d}}$ and between-class covariance matrix $\Sigma_{B}\in\mathrm{R^{d\times d}}$ are:
$\begin{split}\Sigma_{W}:=&\frac{1}{n}\sum_{k=1}^{K}\sum_{j=1}^{n_{k}}(f_{k,j}-\bar{f_{k}})(f_{k,j}-\bar{f_{k}})^{\top}\\ \Sigma_{b}:=&\frac{1}{K}\sum_{k=1}^{K}(\bar{f_{k}}-f_{G})(\bar{f_{k}}-f_{G})^{\top}\\ \end{split}$ (11)
The $\Sigma_{W}$ matrix shows how distant are individual features $f_{k,j}$ from their class prototype $\bar{f_{k}}$ and it is an indicator of feature compactness. The $\Sigma_{b}$ matrix shows how distant are the class prototypes from the global feature, indicating the class separability. Using these matrices we measure the Neural collapse Variability $NC1$ according to [119] as follows:
$NC1:=\frac{1}{K}trace(\Sigma_{W}\Sigma_{b}^{\dagger})$ (12)
where the $\dagger$ symbol denotes the pseudo inverse of $\Sigma_{b}$. $NC1$ measures the magnitude of the within-class covariance $\Sigma_{W}$ compared to the magnitude of the between-class covariance $\Sigma_{b}$ as explained in [119].
In practise, a low $NC1$ measure shows that the model has more compact features since $\Sigma_{W}\downarrow$ decreases and more separable class prototypes because the $\Sigma_{b}\uparrow$ increases. Having more compact features and more separable class prototypes make the representations better and enhance the classification results as shown empirically in previous works [112, 102, 103, 99].
Using Equation 12, we measure the $NC1$ of the deep features of the penultimate layer of SE and APA*, in Table 7 using ImageNet-LT test-set. As the results suggest, our APA* has lower $NC1$ measure for all backbones, showing that APA* produces superior representations that are more compact and seperable than the baseline. This provides another qualitative explanation why our APA* has better performance than SE.
Table 7: Neural Collapse NC1 measure, on ImageNet-LT test set. APA* has lower $NC1$ measure than the baseline, which indicates that it has learned superior representations.
Backbone SE-$NC1\downarrow$ APA*-$NC1\downarrow$
ResNet-50 3.04 2.71
ResNeXt-50 3.38 2.55
ResNet-101 3.15 2.69
ResNet-152 3.24 2.69
Appendix 0.B Implementation Details
The implementation details of APA* and AGLU are shown in Table 8. For balanced ImageNet-1K, the $\lambda$ parameters are initialised as random variables drawn from a Uniform distribution (U), with low parameter 00, and high parameter $1.0$. The APA $\kappa$ parameters are initialised with $U(0,1)$ and the AGLU $\kappa$ parameters are with initialised with $U(0.8,1.2)$. For all other downstream tasks, that use a pretrained model, such as COCO, LVIS, Places-LT and V3Det, we don’t re-initialise the $\kappa$ and $\lambda$ parameters and we simply load them from the pretained ImageNet1K model.
0.B.1 Stable APA implementation
During the development of APA, we found that it is more stable to use Softplus $s_{f}(z,\beta)=\frac{1}{\beta}\ln(1+\exp(\beta z))$, than double exponents, when computing APA. Thus our stable code implementation is:
$\eta_{ad}(z,\kappa,\lambda)=\exp(\frac{1}{\lambda}s_{f}(\kappa z-\ln(\lambda),-1))$ (13)
and it is equivalent to the APA used in the main paper.
Method ImageNet-LT i-Naturalist18 Places-LT C100-LT LVISv1
R50/X50 R50 R152 R32 MRCNN-R50
Batch size 256 1024 256 512 16
Optimiser SGD SGD SGD SGD SGD
LR 0.2 0.5 0.1 0.2 0.02
epochs 200 500 40 500 24
Weight Decay 1e-4 1e-4 5e-5 1e-3 1e-4
Norm Weight Decay 1e-4 0.0 5e-5 1e-3 1e-4
Bias Weight Decay 0.0 0.0 0.0 0.0 0.0
Attention Dropout 0.1 0.0 0.1 0.1 0.1
Mixup $\alpha$ 0.2 0.2 0.2 0.2 -
CutMix $\alpha$ 1.0 1.0 - -
Label smoothing $\epsilon$ 0.1 0.1 - -
Repeated Aug - - - -
AutoAugment - -
RandAugment - - - -
Erasing prob - 0.1 - - -
Cutout - - - -
Cos. Cls. scale 16 16 learnable learnable N/A
Norm. Mask scale N/A N/A N/A N/A learnable
Sampler random random random random RFS
APA $\kappa$ Init U(-1,0) U(0,1) N/A U(-1,0) N/A
APA $\lambda$ Init U(0,1) U(0,1) N/A U(0,1) N/A
AGLU $\kappa$ Init U(0.8,1.2) U(0.8,1.2) N/A U(0.8,1.2) N/A
AGLU $\lambda$ Init U(0,1) U(0,1) N/A U(0,1) N/A
Table 8: Implementation details for Long-tailed Datasets, across various architectures.
0.B.2 AGLU derivatives
Proof of Eq. 9. Then the gradient of AGLU with respect to $\kappa$ is:
$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial\kappa}$ $=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial\kappa}$ (14)
$=x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\lambda x\exp(-\kappa x))$
$=x^{2}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{-1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
$=x^{2}\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
Proof of Eq. 10. Then the gradient of AGLU with respect to $\lambda$ is:
$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial\lambda}$ $=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial\lambda}$ (15)
$=x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(\exp(-\kappa x))$
$=\frac{-x}{\lambda}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
$=\frac{-x}{\lambda}\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
Proof of Eq. 11. Then the gradient of AGLU with respect to $\lambda$ is:
$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial x}$ $=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$ (16)
$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\cdot\partial\frac{(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\kappa\lambda\exp(-\kappa x))$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
Appendix 0.C Results
0.C.1 Experiments with AGLU and plain ResNets
In Table 9, we show the result of AGLU when it applied inside plain ResNet50 models, (i.e. without channel attention). As the Table suggests, by simply replancing the RELU with AGLU, our method consistently increases the performance of plain ResNet models.
Table 9: Top-1 accuracy on long-tailed classification datasets using ResNets.
Dataset CIFAR100-LT ImageNet-LT iNaturalist
Imbalance factor 10 100 256 500
Model ResNet-32 ResNet50 ResNeXt50 ResNet50
RELU 65.7 51.8 55.0 57.0 69.9
AGLU (ours) 66.8 52.0 56.0 57.6 72.4
0.C.2 Experiments with AGLU and Vision Transformers on ImageNet1K
We perform a preliminary experiment with Vision Transformer models such as ViT [19] and Swin [56] using ImageNet1K. We replace the GELU activation with AGLU in every feedforward layer and we keep all other settings the same. As shown in Table 10, AGLU performs comparably to GELU. We believe this is because the Self-Attention function makes the features smooth, by removing their harmonising components and it makes them more Gaussian-like [90, 114, 18, 5]. Consequently, the Gaussial linear error unit, GELU, might be a good choice for the ViT network and our AGLU method has comparable performance.
Table 10: Results of AGLU using ViT models on ImageNet1K.
Model Activation epochs top-1
ViT-B [19] GELU 200 78.3
ViT-B [19] AGLU 200 78.5
Swin-T [56] GELU 100 78.7
Swin-T [56] AGLU 100 78.9
0.C.3 Impact of Initialisation
In all of our experiments, we have initialised $\lambda$ using the Uniform distribution with low parameter 0 and high parameter 1 as a default. Regarding the $\kappa$ parameter inside AGLU, we initialise it to be close to $1.0$, as this works best, as shown in Table 11. Regarding the $\kappa$ parameter inside the attention layer, we found that initialising it with $Uniform(-1,0)$ is slightly better than $Uniform(0,1.0)$ as shown in Table 12.
Table 11: AGLU-$\kappa$ parameter initialisation, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
AGLU - $\kappa$ top-1
$Uniform(0,1)$ 57.7
$Uniform(-2,0)$ 57.3
$Uniform(-3,0)$ 57.4
$Uniform(-2,2)$ Failed
$Uniform(0.8,1.2)$ 57.9
Table 12: $\kappa$ parameter initialisation inside the attention layer, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
APA - $\kappa$ top-1
$Uniform(0,1)$ 57.6
$Uniform(-1,0)$ 57.9
0.C.4 Channel specific $\lambda$ and $\kappa$
We have also tried a variant that uses seperate $\lambda$ and $\kappa$ parameters for every channel. As shown in Table 13, this variant performs worse than using shared $\lambda$ and $\kappa$ parameters for the channels.
Table 13: Results with Channel Specific $\lambda$ and $\kappa$, using APA* ResNet50 backbone on ImageNet-LT.
APA top-1
Channel Specific 57.5
Channel Shared 57.9
0.C.5 Baseline enhancements
We show the detailed ablation study for ImageNet-LT in Table 14. First, the vanilla ResNet50 model trained for 100 epochs on ImageNet-LT achieves $44.4\%$. When we train for 200 epochs then it adds $1.5$pp and switching from linear classifier to cosine classifier adds another $0.4$pp. Stronger training techniques such Mixup [106], Auto-Augment [12] and weight decay tuning further boost the performance by $3.3$pp. Post-calibrated Softmax [34] adds an additional $5.4$pp and finally the Squeeze and Excite module [37] adds another $1.0$pp reaching the final $56.0\%$. Most baseline performance comes from the PC-Softmax and the weight decay finetuning. On top of this strong baseline, our APA increases the performance by $1.0$pp, showing its strong generalisability. Dropout and LayerNorm further increase the performance by $0.4$pp and finally AGLU adds a respectable $0.5$pp reaching $57.9\%$ accuracy on ImageNet-LT. The absolute improvement of all modules is $13.5$pp and our proposed methods, APA and AGLU, contribute by $1.5$pp which is a relative $11\%$ of the total absolute improvement.
Table 14: Detailed Ablation Study, using ResNet50 on ImageNet-LT.
200 epochs Cosine Classifier SE-nets[37] AutoAugment[12] Mixup [106] Weight Decay Tuning [3] PCS [34] APA (ours) Dropout [32] LayerNorm [4] AGLU (ours) ImagetNet-LT
44.4
45.9
46.3
46.8
45.2
45.9
46.6
46.6
49.6
55.0
51.7
56.0
57.0
57.3
57.4
57.9
0.C.6 Qualitative Results
In Figure 6, we show the learned parameters, when training with the balanced and imbalanced ImageNet. Regarding the $\lambda$ inside the AGLU layers in (a), we see that both balanced-trained and imbalanced-trained networks prefer an all-pass filter for the early 2-3 layers, possibly, because the networks are uncertain which features to remove. Then in the intermediate layers, we observe smaller $\lambda$ that corresponds to harder filters and in the final semantic layers we observe larger $\lambda$, possibly, because the network prefers smoother semantic features in order to have smoother classification boundaries. In (b), we see a ‘down-down-up’ $\kappa$-pattern in most bottlenecks, for both balanced and imbalanced ImageNet, showing that the networks prefer softer activations, at first, and harder activations before the residual connection. This indicates that the networks, first, keep most information inside the bottleneck’s projections, and second, they disregard any redundant information, using harder activation, only before performing addition using the residual connection.
Finally, in the last bottlenecks, i.e. layers 45-50, the $\kappa$ parameter diminishes, showing that the network prefers overly smooth activations, possibly, to enhance the classification using smoother classification boundaries.
Regarding the attention layers, the $\kappa$ parameter in (d) dominates over the influence of $\lambda$ in (c), showing that hard channel attention is more preferable than soft channel attention for all layers.
Figure 6: Visualisations of the learned $\lambda$ and $\kappa$ parameters for balanced ImageNet1K (IN-IK) training in blue, and imbalanced ImageNet-LT (IN-LT) training in blue.
0.C.6.1 Visualisations on Imagenet-LT
We further show more qualitative results on ImageNet-LT with ResNet50 backbone. On the left subfigure, we show the model’s highest prediction marked with F,C,R that stands for frequent, common and rare class respectively and the Grad-cam activation [75]. On the right subfigure, we show the last layer’s channel attention signal and its corresponding entropy denoted with (E). As the Figure shows, APA* produces higher entropy attention signals than the baseline and predicts both frequent and rare classes correctly.
Figure 7: Comparative Results between the SE-ResNet50 (baseline) and APA*-ResNet50 (ours) with respect to the activations (left) and the attention entropy (right). F,C,R denote frequent, common and rare samples from ImageNet-LT. Our method produces attention signals that have significantly larger entropy than the baseline for both frequent and rare classes.
0.C.7 Calibration results
Calibration is an important property of models, since it reassures that the confidence of the prediction matches the actual accuracy. When models are not calibrated, then they give wrong predictions with high confidence score (overconfident models) or make correct predictions with low confidence score (underconfident models). In both situations, the miscalibrated models cannot help in the decision making process because their predictions do not reflect their actual accuracy.
In practice in long-tailed learning, the use of complex augmentations and regularisations like mixup, cutmix, label-smoothing, auto-augment and cosine classifier may improve the accuracy but it also reduces the confidence of the model due to over regularisation. As shown in Figure 8 (left-subfigure), SE-Resnet50 is under-confident due to the usage of complex training that includes heavy augmentations and regularisations. When APA* is applied, it reduces the Expected Calibration Error (ECE) as shown in Figure 8 (right-subfigure) for all backbones.
Figure 8: Calibration results using ResNets on ImageNet-LT. SE (left) is underconfident, i.e., its confidence scores are lower than its actually accuracy due to over-regularisation. Our APA* (right) reduces the ECE and makes more accurate predictions with higher confidence than SE.
0.C.8 Next textual token prediction experiment
We perform one preliminary next-token prediction experiment using GPT2 [70] and the FineWeb-Edu [59] subset that contains 10 billion GPT2 tokens. The model is based on the GPT2 smallest architecture, which contains 117M parameters, and the code implementation follows [45]. We train the GPT2 model for one epoch, using 8 V100 GPUs, a total batch size of 0.5M tokens, learning rate $46e-$ 4 and Adam optimizer [48] with momentum. We test the model on the HellaSwag benchmark [105] using zero-shot evaluation. To apply AGLU with GPT2, we simply switch the GELU activation with AGLU inside all MLP layers of the transformer. As the results suggest in Table 15, our AGLU increases the performance of GPT2 by $0.4\%$, showing that AGLU could be a good alternative for text-classification.
Table 15: Comparative results using GPT2 smallest model and HellaSwag benchmark.
Method Accuracy
GELU 31.0
AGLU 31.4

### v3

@@
-Adaptive Parametric Activation
+Adaptive Parametric Activation: Unifying and Generalising Activation Functions Across Tasks
 Abstract
@@
 To this end, we propose the Adaptive Parametric Activation (APA) function, a novel and versatile activation function that unifies most common activation functions under a single formula.
-APA can be applied in both intermediate layers and attention layers, significantly outperforming the state-of-the-art on several imbalanced benchmarks such as ImageNet-LT, iNaturalist2018, Places-LT, CIFAR100-LT and LVIS and balanced benchmarks such as ImageNet1K, COCO and V3DET.
+APA can be applied in both intermediate layers and attention layers, significantly outperforming the state-of-the-art on several imbalanced benchmarks such as ImageNet-LT, iNaturalist2018, Places-LT, CIFAR100-LT and LVIS.
+Also, we extend APA to a plethora of other tasks such as classification, detection, visual instruction following tasks, image generation and next-text-token prediction benchmarks.
+APA increases the performance in multiple benchmarks across various model architectures.
 The code is available at https://github.com/kostas1515/AGLU.
-Keywords:
-Activation function Long-tailed learning
-1 Introduction
-Image recognition has witnessed tremendous progress over the last years due to the use of deep learning, large image datasets such as ImageNet1K [17] and advances in model architectures [19, 28], learning algorithms [24, 69, 25], activation layers [31, 21, 27] and normalisation techniques [4, 40].
+I Introduction
+Image recognition has witnessed tremendous progress over the last years due to the use of deep learning, large image datasets such as ImageNet1K [deng2009imagenet] and advances in model architectures [dosovitskiy2021an, he2016deep], learning algorithms [he2022masked, radford2021learning, he2020momentum], activation layers [hendrycks2016gaussian, glorot2011deep, he2015delving] and normalisation techniques [ba2016layer, ioffe2015batch].
 In this work, we focus on the activation layer of the network.
-In balanced image classification works, it was empirically shown that if the activation function is close to the real data distribution then the model converges faster because the learning objective becomes easier [27, 31].
-Based on this, the GELU [31] and the PRELU [27] were proposed as alternatives to the commonly used RELU [21] and they were utilised inside the model’s layers to activate the intermediate features of the network.
-Similarly, in imbalanced image classification, many works have empirically shown that the Sigmoid or the Softmax activation functions, are inappropriate and using another activation function increases the performance [71, 34, 63, 1].
-Based on that, the Gumbel activation [1] and the Balanced Softmax [71] activation were proposed and they were used inside the classification layer to predict the classes.
+In balanced image classification works, it was empirically shown that if the activation function is close to the real data distribution then the model converges faster because the learning objective becomes easier [he2015delving, hendrycks2016gaussian].
+Based on this, the GELU [hendrycks2016gaussian] and the PRELU [he2015delving] were proposed as alternatives to the commonly used RELU [glorot2011deep] and they were utilised inside the model’s layers to activate the intermediate features of the network.
+Similarly, in imbalanced image classification, many works have empirically shown that the Sigmoid or the Softmax activation functions, are inappropriate and using another activation function increases the performance [Ren2020balms, hong2021disentangling, menon2021longtail, alexandridis2022long].
+Based on that, the Gumbel activation [alexandridis2022long] and the Balanced Softmax [Ren2020balms] activation were proposed and they were used inside the classification layer to predict the classes.
 In contrast to balanced classification, these works focused only on the classification layer and they disregarded the importance of the intermediate activations.
@@
 the classification layer, and the intermediate layers.
+Figure 1: Our APA unifies most activation functions under the same formula.
 Our findings show that the classification logit distribution of a pretrained model heavily depends on the degree of data imbalance.
@@
 To this end, we empirically show that the commonly used Sigmoid activation function cannot generalise for both balanced and imbalanced learning, because it is non-parametric and does not align with the imbalanced data distribution.
-Motivated by this, we develop a novel Adaptive Parametric Activation (APA) function.
+Motivated by this, we develop a novel Adaptive Parametric Activation (APA) function shown in Figure 1.
 APA allows the model to align its activations to both the balanced and imbalanced data distributions and reach great performance in both tasks.
-APA has several benefits, it unifies most previous activations functions such as the Sigmoid, the Gumbel [1], the RELU [21], the SiLU [31] and the GELU[31] under a common formula.
+APA has several benefits, it unifies most previous activations functions such as the Sigmoid, the Gumbel [alexandridis2022long], the RELU [glorot2011deep], the SiLU [hendrycks2016gaussian] and the GELU[hendrycks2016gaussian] under a common formula.
 Also, it uses two learnable parameters that allow the network to select the best activation function during optimisation enlarging the model’s capacity.
 Moreover, our APA is versatile, it can be used as a direct replacement to RELU, or it can replace the Sigmoid activation function inside the attention mechanism boosting the performance significantly and consistently.
-Finally, APA can be generalised to both imbalanced and balanced classification and detection tasks and surpass the state-of-the-art (SOTA).
+Finally, APA can be generalised to both imbalanced and balanced tasks and surpass the state-of-the-art (SOTA).
 Our contributions are:
@@
   • We propose the novel APA function that unifies most common activation functions under a single formula;
-  • We have validated the efficacy of APA on a range of long-tailed benchmarks, including ImageNet-LT [58], iNaturalist18 [84], Places-LT [58], CIFAR100-LT [8], LVIS [22] and balanced benchmarks such as ImageNet1k [17], COCO [55] and V3Det [88] largely surpassing the state of the art.
-2 Preliminaries
-Activation function.
-First, we show the importance of the activation function in model optimisation, following [1].
+  • We have validated the efficacy of APA on a range of long-tailed benchmarks, including ImageNet-LT [liu2019large], iNaturalist18 [van2018inaturalist], Places-LT [liu2019large], CIFAR100-LT [cao2019learning], LVIS [gupta2019lvis] largely surpassing the state of the art.
+  • We comprehensively evaluate APA across various architectures and benchmarks, including image classification, object detection, VQA tasks, image generation and text classification and show that APA generalises to both large scale datasets and foundation models.
+II Related Work
+II-A Long-tailed image recognition.
+Long-tailed image recognition can be grouped according to representation learning and classifier learning techniques.
+Representation learning techniques improve the feature quality through rare class features generation [vigneswaran2021feature, wang2021rsg, zang2021fasa], contrastive objectives [wang2021contrastive, cui2021parametric, zhu2022balanced, samuel2021distributional, li2022targeted, kang2021exploring], ensemble or fusion models [wang2021long, li2022nested, 9774921, zhou2020bbn, li2022trustworthy, cui2022reslt, cai2021ace], knowledge distillation [li2022nested, iscen2021cbd, he2021distilling, li2021self], knowledge transfer [liu2019large, parisot2022long, zhu2020inflated] and data augmentation [zhong2021improving, park2022majority, xu2021towards].
+Classifier learning techniques enhance rare class classification through decoupled training [kang2019decoupling, zhang2021distribution, wang2020devil, kim2020adjusting, hsu2023abc], margin adjustment [menon2021longtail, Ren2020balms, hong2021disentangling, cao2019learning, hyun2022long, pan2021model, wang2021seesaw, wang2022c2am, zhao2022adaptive, alexandridis2023inverse, alexandridis2025fractal], cost-sensitive learning [cui2019class, khan2017cost, wang2017learning, huang2016learning], resampling [chawla2002smote, mahajan2018exploring, shen2016relay, zou2018unsupervised, gupta2019lvis, park2022majority, hong2022safa, zang2021fasa], dynamic loss adaptation based on batch statistics [hsieh2021droploss, tan2020equalization, wang2021adaptive], gradient statistics [tan2021equalization, li2022equalized], weight norms [wang2022c2am] and classification scores [feng2021exploring, he2022relieving].
+These techniques are efficient, however they are applied either during the input phase or during the loss disregarding the intermediate activations.
+II-B Attention Networks.
+Attention networks, such as spatial attention [jaderberg2015spatial], channel attention [hu2018squeeze] and spatial-channel attention [woo2018cbam], are widely used in image recognition.
+Self-attention is the core mechanism of the Visual Transformer (ViT) architecture [dosovitskiy2021an], and recently many improvements have been made to the Transformer architecture [touvron2021training, touvron2021going, liu2021swin, xiao2021early] and its training procedure [steiner2021train, beyer2022better, touvron2022deit].
+APA can be combined with many attention models like channel attention, Spatial, Spatial-Channel attention and self-attention models and further boost the performance.
+II-C Activation functions
+The RELU function [glorot2011deep] dominates the landscape of deep image classification, and it is especially used inside the convolutional networks, while the GELU and SiLU [hendrycks2016gaussian] are more commonly used inside the transformer network [dosovitskiy2021an] or the ConNext models [liu2022convnet, woo2023convnext].
+The Gumbel activation is effective for imbalanced image recognition tasks [alexandridis2022long], and recently, it showed good performance for balanced tasks as well [das2025gompertz].
+The PRELU function [he2015delving] is a generalisation of the RELU, because it linearly activates both positive and negative inputs and the ELU [clevert2015fast] is a follow-up work that enforces saturation on the negative inputs after some threshold.
+Compared to those, our APA, generalises the RELU, the GELU, SiLU and Gumbel and it can also be seen as a smoother version of PRELU.
+Recently, there has been increased interest in diffeomorphic activation functions used in image classification [chelly2024trainable], segmentation [freifeld2017transformations], parameter-efficient tuning [mantri2024digraf] and graph neural networks [mantri2024digraf].
+However, these activation functions require extensive forward and backward computations and a user defined input-domain interval.
+In contrast, APA forward and backward is straightforward and it operates in the whole input domain.
+III Preliminaries
+III-A Activation function.
+First, we show the importance of the activation function in model optimisation, following [alexandridis2022long].
 Let’s consider the example of binary classification, where $z$ is the input, $y\in\{0,1\}$ is the target class and $f(z)=W^{T}z+b$ is the classification network.
@@
 The probability of the class $P(y=1)$ is:
-$P(y=1)=P(f(z)+\epsilon>0)=P(\epsilon>-f(z))=1-F(-f(z))$ (2)
+$P(y=1)=P(f(z)+\epsilon>0)=P(\epsilon>-f(z))$ (2)
+$=1-F(-f(z))$
 where $F$ is the cumulative distribution function.
 During network optimisation, if one chooses the Sigmoid activation function $\sigma(z)=(1+\exp(-z))^{-1}$, then the prediction $\bar{y}$ is obtained using $\bar{y}=\sigma(f(z))$ and the probability $P(\bar{y}=1)$ is:
-$P(\bar{y}=1)=\frac{1}{1+\exp(-f(z))}=F_{\text{logistic}}(f(z);0,1)=1-F_{\text{logistic}}(-(f(z);0,1)$ (3)
+$P(\bar{y}=1)=\frac{1}{1+\exp(-f(z))}=F_{\text{logistic}}(f(z);0,1)$ (3)
+$=1-F_{\text{logistic}}(-(f(z);0,1)$
 Comparing Eq.
@@
 3, it is shown that when we use the Sigmoid, we assume that the error term $\epsilon$ follows the standard Logistic distribution.
-If we use Gumbel activation then we assume that the error follows the Gumbel distribution [1] and, in general, any activation function assumes a different distribution of $\epsilon$.
+If we use Gumbel activation then we assume that the error follows the Gumbel distribution [alexandridis2022long] and, in general, any activation function assumes a different distribution of $\epsilon$.
 For this reason, the activation function can be seen as an initialisation point, or a prior belief of how the real data are distributed.
-If the prior is good, then the learning objective becomes easier and the performance is increased as it was shown empirically by many past studies [27, 1, 21, 77].
-2.1 Balanced versus Imbalanced learning
+If the prior is good, then the learning objective becomes easier and the performance is increased as it was shown empirically by many past studies [he2015delving, alexandridis2022long, glorot2011deep, skorski2021revisiting].
+III-B Balanced versus Imbalanced learning
 In this subsection, we perform a statistical analysis on the empirical distributions of the logits and the intermediate activations, to understand the importance of the activation function in balanced and imbalanced learning.
 Classification logits.
-We train a MaskRCNN [26] ResNet50[28] on LVIS dataset [22], which is a highly imbalanced object detection dataset containing 1203 classes.
+We train a MaskRCNN [he2017mask] ResNet50[he2016deep] on LVIS dataset [gupta2019lvis], which is a highly imbalanced object detection dataset containing 1203 classes.
 After the model has converged, we perform inference and store the predicted classification logits, i.e., $f(z_{y})$ for every class $y$.
@@
 Regarding its distribution, it resembles the Gumbel distribution, because it has a heavy right tail and it is skewed.
-To quantify that, we calculate the statistical distance, using the Kolmogorov-Smirnov (KS) test [62], between all empirical class distributions and the Gumbel and Logistic theoretical distributions, shown in (a) and (e).
+To quantify that, we calculate the statistical distance, using the Kolmogorov-Smirnov (KS) test [massey1951kolmogorov], between all empirical class distributions and the Gumbel and Logistic theoretical distributions, shown in (a) and (e).
 As shown in (d), most classes have smaller distance to the Gumbel distribution.
@@
 In general, the logit distributions are closer to the logistic distribution as shown in (h).
-This explains why the Sigmoid activation achieves better performance in the balanced classification task and worse performance in imbalanced classification [1, 71, 34, 63].
+This explains why the Sigmoid activation achieves better performance in the balanced classification task and worse performance in imbalanced classification [alexandridis2022long, Ren2020balms, hong2021disentangling, menon2021longtail].
 Next, we show that data imbalance also affects the intermediate layers, by studying the channel attention as an example.
@@
 Attention mechanism for input $X\in\mathbb{R}^{H\times W\times C}$ re-weights the input features $X$ by applying an attention function $A(X)$, i.e., $X^{\prime}=A(X)\otimes X$.
-For example, in Channel Attention (CA) [37], $X^{\prime}=\sigma(\text{MLP}(\text{GAP}_{c}(X)))\otimes X$, where $\text{GAP}_{c}\in\mathbb{R}^{1\times 1\times C}$ is Global Average Pooling and $\otimes$ is the element-wise product.
+For example, in Channel Attention (CA) [hu2018squeeze], $X^{\prime}=\sigma(\text{MLP}(\text{GAP}_{c}(X)))\otimes X$, where $\text{GAP}_{c}\in\mathbb{R}^{1\times 1\times C}$ is Global Average Pooling and $\otimes$ is the element-wise product.
 In this case, the attention function is $A_{CA}(X)=\sigma(\text{MLP}(\text{GAP}_{c}(X)))$.
@@
 In (a) the attention signals when training with imbalanced ImageNet and balanced ImageNet have similar variance in the first layer but completely different in the last most semantic layer in (b).
-The variance with ImageNet-LT training drops to zero for deeper layers as shown in (c), because the attention promotes only few frequent classes.
+The variance with ImageNet-LT training drops to zero for deeper layers as shown in (c), because the attention promotes only the frequent classes.
 In (d) and (e) the entropy of channel attention is smaller for the rare classes than the frequent classes in both i-Naturalist18 and ImageNet-LT training.
-In (f) and (g) the channel responses are smaller for the rare classes than the frequent ones for i-Naturalist18 and ImageNet-LT respectively.
+In (f) and (g) the channel responses are smaller for the rare classes forboth i-Naturalist18 and ImageNet-LT.
+These observations shows that channel attention is not robust during imbalanced training, motivating the development of APA.
 Balanced vs Imbalanced channel attention.
-We train SE-ResNet50 [37] models (SE-R5) on balanced ImageNet-1K and imbalanced ImageNet-LT.
+We train SE-ResNet50 [hu2018squeeze] models (SE-R50) on balanced ImageNet-1K and imbalanced ImageNet-LT.
 After training the models, we analyse the average channel attention signals for a random batch of 128 test images.
 Fig.
-2-a shows the output $A_{CA}$, in the first layer of SE-R50, Fig.
-2-b shows the output, in the last layer, and Fig.
-2-c shows the variance of $A_{CA}$ across all layers.
+3-a shows the output $A_{CA}$, in the first layer of SE-R50, Fig.
+3-b shows the output, in the last layer, and Fig.
+3-c shows the variance of $A_{CA}$ across all layers.
 As shown in Fig.
-2-a, the attention is similar in the first layer and it is different in the last layer, as shown in Fig.
-2-b.
+3-a, the attention is similar in the first layer and it is different in the last layer, as shown in Fig.
+3-b.
 As displayed in Fig.
-2-c, the attention variance with ImageNet1K (blue-curve) is larger and the attention reweights all channels and affects all classes.
+3-c, the attention variance with ImageNet1K (blue-curve) is larger and the attention reweights all channels and affects all classes.
 In contrast, for the imbalanced case (orange-curve), the attention signal has small variance, indicating that it is biased to some classes.
@@
 Since the channel attention produces a probabilistic weighting vector via the Sigmoid activation, we calculate the total entropy of channel attention, for a layer $l$, as the sum of the binary channel distribution entropies as follows:
-$\begin{split}E_{l}=-\frac{1}{C}\sum_{i=1}^{C}[(A_{CA}(X_{l})\log(A_{CA}(X_{l}))+(1-A_{CA}(X_{l})\log(1-A_{CA}(X_{l}))]\end{split}$ (4)
+$\begin{split}E_{l}=-\frac{1}{C}\sum_{i=1}^{C}[(A_{CA}(X_{l})\log(A_{CA}(X_{l}))\\ +(1-A_{CA}(X_{l})\log(1-A_{CA}(X_{l}))]\end{split}$ (4)
 When the layer’s entropy is closer to zero, the channel attention signals are closer to 1 and they do not affect the original features for that layer, i.e., $X^{\prime}=1\otimes X$.
@@
 In conclusion, this analysis shows that data imbalance affects the quality of the activations inside the intermediate layers and it highlights the limitation of the Sigmoid activation to model the rare classes.
-Figure 3: Our APA unifies most activation functions under the same formula.
-3 Method
+IV Method
+IV-A Adaptive Parametric Activation
 As shown in the previous section, the degree of data imbalance affects both the classification logits and the intermediate layers.
-While it is possible to perform a statistical analysis and select the appropriate activation for the classification layer, this is difficult to do for all layers of the network, because first, the data distributions inside the layers dynamically change during training  [4, 40], and secondly, there is no one-to-one correspondence between classes and intermediate channels, which hinders attribution.
+While it is possible to perform a statistical analysis and select the appropriate activation for the classification layer, this is difficult to do for all layers of the network, because first, the data distributions inside the layers dynamically change during training  [ba2016layer, ioffe2015batch], and secondly, there is no one-to-one correspondence between classes and intermediate channels, which hinders attribution.
 For this reason, we propose the Adaptive Parametric Activation (APA):
@@
 $\lambda\in(0,\infty)$ is the asymmetrical parameter that controls the function’s response rate to positive and negative inputs, allowing the model to have different learning degrees when the input is positive or negative.
-This function is also known as Richard’s curve [73] and it unifies the most common activation functions.
+This function is also known as Richard’s curve [richards1959flexible] and it unifies the most common activation functions.
 For example, if $\kappa=\lambda=1$ then APA becomes the Sigmoid activation that has a symmetric response rate for both positive and negative inputs and it is successful for balanced classification tasks.
-If $\kappa=1,\lambda\xrightarrow[]{}0$ then APA becomes the Gumbel activation that has an asymmetric response rate and it is successful for long-tailed instance segmentation tasks [1].
+If $\kappa=1,\lambda\xrightarrow[]{}0$ then APA becomes the Gumbel activation that has an asymmetric response rate and it is successful for long-tailed instance segmentation tasks [alexandridis2022long].
 This behaviour is shown in Fig.
@@
 5, we also define the Adaptive Generalised Linear Unit (AGLU):
-$\begin{split}AGLU(z,\kappa,\lambda)=z\cdot\eta_{ad}(z,\kappa,\lambda)\\ \end{split}$ (6)
+[…]
 Figure 4: AGLU derivatives with respect to $\kappa$ (top), $\lambda$ (middle) and $z$ (bottom).
-AGLU has many interesting properties, for example, if $\kappa=\lambda=1$ then AGLU becomes the Sigmoid Linear Unit (SiLU) [31].
-If $\kappa=1.702,\lambda=1$, then it becomes the Gaussian Error Linear Unit (GELU) [31].
-If $\kappa\xrightarrow[]{}\infty$, then AGLU becomes ReLU[21] and if $\lambda\xrightarrow[]{}\infty$, then AGLU becomes the identity function, as shown in Fig.
-3-right.
+AGLU has many interesting properties, for example, if $\kappa=\lambda=1$ then AGLU becomes the Sigmoid Linear Unit (SiLU) [hendrycks2016gaussian].
+If $\kappa=1.702,\lambda=1$, then it becomes the Gaussian Error Linear Unit (GELU) [hendrycks2016gaussian].
+If $\kappa\xrightarrow$$\infty$, then AGLU becomes ReLU[glorot2011deep] and if $\lambda\xrightarrow$$\infty$, then AGLU becomes the identity function, as shown in Fig.
+1-right.
 In other words, the $\kappa$ parameter controls the RELU-ness and the $\lambda$ parameter controls the leakage.
 Also, AGLU could be seen as a smoother version of PRELU, because $AGLU(z,1,\lambda\xrightarrow[]{}\infty)=PRELU(z,1)$.
-We compare $AGLU$ to most common activation functions, in more detail, in Table 1.
+We compare AGLU to most common activation functions, in more detail, in Table I.
 The derivative of AGLU with respect to $\kappa$ is:
@@
 Using various $\kappa$ and $\lambda$ combinations AGLU has drastically different behaviour and this enhances the capacity of the network and achieves good performance as shown in the experiments.
-Table 1: Comparison of different activation functions.
-Name Formula Range
-RELU [21] $\eta(z)=\max(0,z)$ $(0,\infty)$
-Gaussian Unit [31] $\eta(z)=z\sigma(1.702z)$ $(-0.17,\infty)$
-Sigmoid Unit [31] $\eta(z)=z\sigma(z)$ $(-0.28,\infty)$
-Mish [64] $\eta(z)=z\tanh(\ln(1+\exp(z))$ $(-0.31,\infty)$
-PRELU [27] $\eta(z,\kappa)=\max(0,z)+\kappa\min(0,z)$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
-ELU [11] $\eta(z,\kappa)=\max(0,z)+\kappa(\exp(\min(0,z))-1)$ $(-\kappa,\infty)$
-AGLU (ours) $\eta(z,\kappa,\lambda)=z\cdot(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
-In the end, our APA is versatile because it can be used not only as an error unit, replacing RELU, but also as an activation function inside the attention mechanism, replacing the Sigmoid activation.
-4 Experiment Setup
-Datasets.
-We use CIFAR100-LT [8] with exponential imbalance factor of 100 and 10, ImageNet-LT [58], Places-LT [58] and iNaturalist2018 [84] following the common long-tailed classification protocol.
-We report our results using top-1 accuracy on the balanced test sets, to fairly evaluate all classes.
-For ImageNet-LT, we split the categories according to their class frequency in the training set, into Many ($>$100 images), Medium (20$\sim$100 images) and Low ($<$20 images) and do per-group evaluation following [58].
-Also, we use the LVISv1 [22] instance segmentation dataset, which has 100K training images and 1203 classes, that are grouped according to Frequent ($>$100 images), Common (10$\sim$100 images) and Rare ($<$10 images) classes.
-For this dataset, we report mask average precision $AP$, bounding box average precision $AP^{b}$ and $AP^{r}$, $AP^{c}$ and $AP^{f}$ which is mask average precision for rare, common and frequent classes respectively.
-Regarding balanced training, we perform experiments on ImageNet1K [17], COCO [55] and the recently proposed V3Det[88], which is a challenging large scale detection dataset with 13K classes and 243K images.
-We report top-1 accuracy for ImageNet1K, $AP^{b}$ and $AP^{m}$ for COCO and $AP^{b}$ for V3Det.
-Table 2: Top-1 accuracy (%) on ImageNet-LT test set. E denotes ensemble.
+TABLE I: Comparison of different activation functions.
+Name Formula Output Range
+RELU [glorot2011deep] $\eta(z)=\max(0,z)$ $(0,\infty)$
+Gaussian Unit [hendrycks2016gaussian] $\eta(z)=z\sigma(1.702z)$ $(-0.17,\infty)$
+Sigmoid Unit [hendrycks2016gaussian] $\eta(z)=z\sigma(z)$ $(-0.28,\infty)$
+Mish [misra2019mish] $\eta(z)=z\tanh(\ln(1+\exp(z))$ $(-0.31,\infty)$
+PRELU [he2015delving] $\eta(z,\kappa)=\max(0,z)+\kappa\min(0,z)$ $(-\infty,\infty)$
+ELU [clevert2015fast] $\eta(z,\kappa)=\max(0,z)+\kappa(\exp(\min(0,z))-1)$ $(-\kappa,\infty)$
+AGLU (ours) $\eta(z,\kappa,\lambda)=z\cdot(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}$ $(-\infty,\infty)$
+APA is versatile and it has many use cases. For example it can be uses in the intermediate layers, i.e. replacing RELU, or inside the attention mechanism, replacing the Sigmoid activation.
+IV-B Non-Linearity and Adaptability
+APA ’s parameters control the amount of non-linearity added to the model.
+To measure this, we follow the newly proposed Neural Redshift framework [teney2024neural] and use the affinity score $\rho_{aff}$ [bouniot2025alexnet] to study the behavior of a randomly initialised three hidden-layer MLP with a scalar output.
+We initialise the MLP weights using various schemes such as Xavier [glorot2010understanding] and Kaiming [he2015delving] and the MLP biases with the uniform distribution U(1,1)U(-1,1), we feed an input of $64^{2}$ evenly spaced 2D coordinates on the [1,1][-1,1] grid and visualise the output in grayscale in Figure 5.
+If RELU is used as nonlinearity, then the MLP has the biggest amount of non-linearity, because it produces the smallest affinity scores $\rho_{aff}$ and the most complex outputs for all the initialisation schemes as shown in the first column.
+If AGLU with small $\lambda$ and larger $\kappa$ values is used then the AGLU-MLP becomes similar to the RELU-MLP, i.e.
+it is more non-linear and it has smaller affinity scores $\rho_{aff}$ and more complex outputs as shown in the the second column.
+Decreasing the $\kappa$ or increasing the $\lambda$ values of AGLU makes the MLP less non-linear, i.e.
+the affinity scores increase and the outputs become smoother as shown in the third and fourth column respectively.
+This has two effects, it allows the users to smoothly control the inductive bias of the model, i.e.
+by fixing the amount of non-linearity or it enables to learn the best nonlinearities by optimization.
+Figure 5: Visualisation of the output of a random MLP, initialised with Xavier-Uniform, Xavier-Normal, Kaiming-Uniform and Kaiming-Normal, across the rows, and $RELU(z)$, $AGLU(z,2,0)$, $AGLU(z,1,0)$ and $AGLU(z,1,1)$ across the columns.
+The RELU MLP is the most non-linear, for all weight initialisations, because it has the smallest affinity scores $\rho_{aff}$ and most complex outputs as shown in the first column.
+Increasing the AGLU’s $\kappa$ value lowers $\rho_{aff}$ and makes the MLP more non-linear, as shown in the second column, while decreasing the $\kappa$ or increasing the $\lambda$ values increases $\rho_{aff}$ and makes it more linear as shown in the third and forth columns respectively.
+V Experiments in Long-tail learning
+V-A Datasets.
+[…]
+TABLE II: Top-1 accuracy (%) on ImageNet-LT test set. E denotes ensemble.
 Method Backbone Many Medium Few Average
-MiSLAS [111] R50 [28] 61.7 51.3 35.8 52.7
-KCL [43] 61.8 49.4 30.9 51.5
-TSC [53] 63.5 49.7 30.4 52.4
-RIDE (3E)+CMO[68] 66.4 53.9 35.6 56.2
-DOC [86] 65.1 52.8 34.2 55.0
-CC-SAM [116] 61.4 49.5 37.1 52.4
-Our Baseline SE-R50 [37] 66.2 53.1 37.1 56.0
+MiSLAS [zhong2021improving] R50 61.7 51.3 35.8 52.7
+KCL [kang2021exploring] 61.8 49.4 30.9 51.5
+TSC [li2022targeted] 63.5 49.7 30.4 52.4
+RIDE (3E)+CMO[park2022majority] 66.4 53.9 35.6 56.2
+DOC [wang2022towards] 65.1 52.8 34.2 55.0
+CC-SAM [zhou2023class] 61.4 49.5 37.1 52.4
+Our Baseline SE-R50 66.2 53.1 37.1 56.0
 APA* (ours) 67.5 54.3 39.3 57.4
-APA* + AGLU (ours) 68.3${}^{\textbf{+1.9}}$ 54.8${}^{\textbf{+0.9}}$ 39.4${}^{\textbf{+2.1}}$ 57.9${}^{\textbf{+1.7}}$
-RIDE (4E) [95] X50 [100] 68.2 53.8 36.0 56.8
-SSD [52] 66.8 53.1 35.4 56.0
-BCL [117] 67.9 54.2 36.6 57.1
-CNT  [67] 63.2 52.1 36.9 54.2
-ALA  [110] 64.1 49.9 34.7 53.3
-ResLT  [13] 63.6 55.7 38.9 56.1
-ABC-Norm [36] 60.7 49.7 33.1 51.7
-RIDE (3E)+CMO+CR [60] 67.3 54.6 38.4 57.4
-LWS+ImbSAM [115] 63.2 53.7 38.3 55.3
-Our Baseline SE-X50 [37] 67.9 53.0 37.7 56.7
+[…]
+RIDE (4E) [wang2021long] X50 68.2 53.8 36.0 56.8
+SSD [li2021self] 66.8 53.1 35.4 56.0
+BCL [zhu2022balanced] 67.9 54.2 36.6 57.1
+CNT  [parisot2022long] 63.2 52.1 36.9 54.2
+ALA  [zhao2022adaptive] 64.1 49.9 34.7 53.3
+ResLT  [cui2022reslt] 63.6 55.7 38.9 56.1
+ABC-Norm [hsu2023abc] 60.7 49.7 33.1 51.7
+RIDE (3E)+CMO+CR [ma2023curvature] 67.3 54.6 38.4 57.4
+LWS+ImbSAM [zhou2023imbsam] 63.2 53.7 38.3 55.3
+Our Baseline SE-X50 67.9 53.0 37.7 56.7
 APA* (ours) 68.9 55.4 39.4 58.4
-APA* + AGLU (ours) 69.8${}^{\textbf{+1.6}}$ 55.70.${}^{\text{0.0}}$ 41.1+2.${}^{\textbf{+2.2}}$ 59.1${}^{\textbf{+1.7}}$
-Implementation Details We primarily use Squeeze and Excite [37] as our baseline with ResNet-32 [28] for CIFAR-LT, ResNet50 for iNaturalist, ResNet50 and ResNext50 [100] for ImageNet-LT and ResNet152 for Places-LT, which has been pretrained on ImageNet1K according to [58].
-For LVIS, we use SE-Resnets with MaskRCNN [26], FPN [54], Normalised Mask [89], RFS sampler [22] and GOL loss [1] as a baseline.
-For V3Det, we use SE-ResNet50 with FasterRCNN [72], FPN and Wrapped Cauchy classifier [23].
-All baselines use bag of tricks [109], and strong training techniques that we discuss and ablate in the Appendix.
-Our motivation for using bag of tricks is two-fold, first, it pushes the performance even further and secondly, it showcases the generalisability of our work.
-For our attention models, we replace the Sigmoid with APA, and we further use LayerNorm [4] and attention dropout [32] with $p=0.1$, for all datasets except iNaturalist.
+[…]
+V-B Implementation Details
+We primarily use Squeeze and Excite [hu2018squeeze] as our baseline with ResNet-32 [he2016deep] for CIFAR-LT, ResNet50 for iNaturalist, ResNet50 and ResNext50 [xie2017aggregated] for ImageNet-LT and ResNet152 for Places-LT, which has been pretrained on ImageNet1K according to [liu2019large].
+For LVIS, we use SE-Resnets with MaskRCNN [he2017mask], FPN [lin2017feature], Normalised Mask [wang2021seesaw], RFS sampler [gupta2019lvis] and GOL loss [alexandridis2022long] as a baseline.
+All baselines use bag of tricks [zhang2021bag], and strong training techniques that we discuss and ablate in the Appendix.
+Our motivation for using bag of tricks is two-fold, first, it pushes the performance even further and secondly, it showcases the efficacy of our work.
+For our attention models, we replace the Sigmoid with APA, and we further use LayerNorm [ba2016layer] and attention dropout [hinton2012improving] with $p=0.1$, for all datasets except iNaturalist.
 We denote this configuration as APA* in our Tables and we ablate its components.
-For our AGLU models, we simply use it in-place of RELU.
-5 Results
-Table 3: (a) Results of APA* and AGLU for iNaturalist and Places-LT. (b) Results of APA* and AGLU for CIFAR100-LT with imbalance 10 and 100.
+For our AGLU models, we simply use it in-place of RELU, or GELU.
+Following [he2015delving] we disable the weight decay of AGLU parameters, allowing the activation curve to reach the optimal shapes during training.
+TABLE III: (a) Results of APA* and AGLU for iNaturalist and Places-LT. (b) Results of APA and AGLU for CIFAR100-LT with imbalance 10 and 100.
 Method iNat18 PlacesLT
-DisAlign [108] 70.6 39.3
-MisLAS [111] 71.6 40.4
-LADE [34] 70.0 38.8
-ALA [110] 70.7 40.1
-TSC [53] 69.7 -
-CNT [67] - 39.2
-WD+MaxNorm [3] 70.2 -
-DOC [86] 71.0 -
-BCL[117] 71.8 -
-ResLT [13] 70.5 39.8
-IIF [2] - 40.2
-ABC Norm [36] 71.4 -
-LWS+ImbSAM[115] 71.1 -
-CC-SAM[116] 70.9 40.6
-AREA [10] 68.4 -
-Our Baseline with SE 71.3 40.5
+DisAlign [zhang2021distribution] 70.6 39.3
+MisLAS [zhong2021improving] 71.6 40.4
+LADE [hong2021disentangling] 70.0 38.8
+ALA [zhao2022adaptive] 70.7 40.1
+TSC [li2022targeted] 69.7 -
+CNT [parisot2022long] - 39.2
+WD+MaxNorm [alshammari2022long] 70.2 -
+DOC [wang2022towards] 71.0 -
+BCL[zhu2022balanced] 71.8 -
+ResLT [cui2022reslt] 70.5 39.8
+IIF [alexandridis2023inverse] - 40.2
+ABC Norm [hsu2023abc] 71.4 -
+LWS+ImbSAM[zhou2023imbsam] 71.1 -
+$C$-SAM[zhou2023class] 70.9 40.6
+AREA [chen2023area] 68.4 -
+Baseline with SE 71.3 40.5
 APA* (ours) 72.3 41.3
-APA* +AGLU (ours) 74.8${}^{\textbf{+3.0}}$ 42.0${}^{\textbf{+1.4}}$
+[…]
 Method 10 100
-BALMS [71] 63.0 50.8
-RIDE (4E) [95] - 49.4
-ACE (4E) [7] - 49.6
-DiVE [29] 62.0 45.4
-SSD [52] 62.3 46.0
-MisLas [111] 63.2 47.0
-HSC [91] 63.1 46.7
-LADE [34] 61.7 45.4
-ResLT (3E) [13] 63.7 49.7
-TLC (4E) [50] - 49.8
-TSC [53] 59.0 43.8
-RIDE+CMO [68] 60.2 50.0
-CC-SAM [116] - 50.8
-RIDE+CMO+CR[60] 61.4 50.7
-AREA [10] 60.8 48.9
-Our Baseline with SE 65.2 50.9
+BALMS [Ren2020balms] 63.0 50.8
+RIDE (4E) [wang2021long] - 49.4
+ACE (4E) [cai2021ace] - 49.6
+DiVE [he2021distilling] 62.0 45.4
+SSD [li2021self] 62.3 46.0
+MisLas [zhong2021improving] 63.2 47.0
+HSC [wang2021contrastive] 63.1 46.7
+LADE [hong2021disentangling] 61.7 45.4
+ResLT (3E) [cui2022reslt] 63.7 49.7
+TLC (4E) [li2022trustworthy] - 49.8
+TSC [li2022targeted] 59.0 43.8
+RIDE+CMO [park2022majority] 60.2 50.0
+CC-SAM [zhou2023class] - 50.8
+RIDE+CMO+CR[ma2023curvature] 61.4 50.7
+AREA [chen2023area] 60.8 48.9
+Baseline with SE 65.2 50.9
 APA* (ours) 65.7 51.9
-APA*+AGLU(ours) 66.8${}^{\textbf{+1.6}}$ 52.3${}^{\textbf{+1.4}}$
-Long-tailed Classification Benchmark.
-We compare APA* against ensemble and fusion models [95, 13, 50, 7], margin adjustment [34, 108, 110], contrastive learning [91, 53, 43], knowledge transfer [67], knowledge distillation [29, 52], decoupled methods [111, 2, 108, 36], sharpness aware minimisation [115, 116, 60] and data augmentation [111, 68].
-On ImageNet-LT, as shown in Table 2, our baseline models with bag of tricks reach the state-of-the-art (SOTA) for both SE-ResNet50 (R50) and SE-ResNeXt50 (X50).
+[…]
+V-C Long-tailed Classification Benchmark.
+We compare APA* against ensemble and fusion models [wang2021long, cui2022reslt, li2022trustworthy, cai2021ace], margin adjustment [hong2021disentangling, zhang2021distribution, zhao2022adaptive], contrastive learning [wang2021contrastive, li2022targeted, kang2021exploring], knowledge transfer [parisot2022long], knowledge distillation [he2021distilling, li2021self], decoupled methods [zhong2021improving, alexandridis2023inverse, zhang2021distribution, hsu2023abc], sharpness aware minimisation [zhou2023imbsam, zhou2023class, ma2023curvature] and data augmentation [zhong2021improving, park2022majority].
+On ImageNet-LT, as shown in Table II, our baseline models with bag of tricks reach the state-of-the-art (SOTA) for both SE-ResNet50 (R50) and SE-ResNeXt50 (X50).
 We want to point out that most performance comes from the bag of tricks and not the SE module, as shown, in detail in the Appendix.
-Our APA* outperforms the SE-R50 baseline by $1.4$ percentage points (pp) on average, by $1.3$pp on frequent categories, $1.2$pp on medium and $2.2$pp on few classes.
+Our APA* outperforms the SE-R50 baseline by $1.4$ percentage points (pp) on average, by 1.31.3pp on frequent categories, 1.21.2pp on medium and 2.22.2pp on few classes.
 Most importantly, it increases the performance of both frequent and rare classes, which is a unique advantage compared to the previous works.
-Additionally, our APA* exceeds RIDE with 3 Experts (3E) and CMO [68] by $1.2$pp on average and by $3.7$pp on the rare classes using a single model.
-When AGLU is combined, it pushes the performance of APA* R50, by $0.5$pp on average, by $0.8$pp on the frequent, $0.5$pp on the medium and $0.1$pp on the few classes.
-APA* with X50 increases the SE-X50 baseline by $1.7$pp on average, $1.0$pp the many classes, $1.4$pp the medium classes and $1.7$pp the few classes.
-Furthermore, it outperforms RIDE(3E)+CMO+CR [60] by $1.0$pp on average, $1.6$pp on frequent classes, $0.8$ on common and $1.0$pp on rare classes using a single model.
-When AGLU is combined to APA* X50, it adds $0.7$pp on average, $0.9$pp on the frequent categories, $0.3$pp on the medium, and $1.7$pp on the few classes.
-Table 4: For all ablations we use ImageNet-LT.
+Additionally, our APA* exceeds RIDE with 3 Experts (3E) and CMO [park2022majority] by 1.21.2pp on average and by 3.73.7pp on the rare classes using a single model.
+When AGLU is combined, it pushes the performance of APA* R50, by 0.50.5pp on average, by 0.80.8pp on the frequent, 0.50.5pp on the medium and 0.10.1pp on the few classes.
+APA* with X50 increases the SE-X50 baseline by 1.71.7pp on average, 1.01.0pp the many classes, 1.41.4pp the medium classes and 1.71.7pp the few classes.
+Furthermore, it outperforms RIDE(3E)+CMO+CR [ma2023curvature] by 1.01.0pp on average, 1.61.6pp on frequent classes, $0.8$ on common and 1.01.0pp on rare classes using a single model.
+When AGLU is combined to APA* X50, it adds 0.70.7pp on average, 0.90.9pp on the frequent categories, 0.30.3pp on the medium, and 1.71.7pp on the few classes.
+TABLE IV: For all ablations we use ImageNet-LT.
 In (a) we compare adaptive activation to other learnable activation functions using SE-ResNet50.
@@
 with Temp 66.9 53.4 39.7 56.7
-[…]
+APA 67.1 53.8 39.6 57.0
 Attention type Avg
-Spatial [98] 54.8
+Spatial [woo2018cbam] 54.8
 +APA 55.2
 +APA + AGLU 56.4
-Spatial + Channel  [98] 55.6
+Spatial + Channel  [woo2018cbam] 55.6
 +APA 56.9
@@
 CE 51.7 52.9
-PC-Softmax [34] 56.0 57.4
-cRT [44] 55.6 56.4
-Decoupled-DRW  [8] 55.1 56.6
-BSCE [71] 56.0 57.0
+PC-Softmax [hong2021disentangling] 56.0 57.4
+cRT [kang2019decoupling] 55.6 56.4
+Decoupled-DRW  [cao2019learning] 55.1 56.6
+BSCE [Ren2020balms] 56.0 57.0
 (a) APA comparison.
@@
 SE-R101 67.5 53.3 37.6 56.7
-[…]
+APA*-R101 68.1 56.0 42.1 58.8
 SE-R152 68.0 54.4 39.8 57.6
-[…]
+APA*-R152 69.0 56.8 41.2 59.4
 SE APA Dropout LayerNorm Avg
@@
 ReLU 57.4
-PReLU [27] 54.8
-ELU [11] 52.6
-Mish [64] 57.4
-GELU [31] 57.5
-SiLU [31] 57.1
+PReLU [he2015delving] 54.8
+ELU [clevert2015fast] 52.6
+Mish [misra2019mish] 57.4
+GELU [hendrycks2016gaussian] 57.5
+SiLU [hendrycks2016gaussian] 57.1
 AGLU 57.9
@@
 (f) AGLU comparison.
-On iNaturalist18, Places-LT and CIFAR100-LT, our SE baselines with bag of tricks again reach the SOTA, as shown in Table 3-a and b.
-Regarding, iNaturalist18, APA* improves the SE-baseline by $1.0$pp.
-It also outperforms BCL [117] by $0.5$pp, ResLT by $1.8$pp and LWS+ImbSAM [115] by $1.2$pp.
-AGLU further enhances the performance by a staggering $3.5$pp compared to SE-baseline, which is a significant increase.
-On Places-LT, our APA* also increases the performance of the SE baseline by $0.8$pp.
-Moreover, it surpasses MisLAS [111] by $0.9$pp and CC-SAM [116] by $0.7$pp.
-When AGLU is combined with APA*, it further increases the performance by $0.7$pp reaching $42.0\%$.
-As shown in Table 3-b, on CIFAR100-LT, APA* improves the performance by $0.5$pp and $1.0$pp compared to SE baseline using an imbalance factor of 10 and 100 respectively.
-Furthermore, compared to RIDE+CMO [68], CC-SAM [116] and RIDE+CMO+CR [60], APA* achieves $1.9$pp, $1.1$pp and $1.2$pp higher accuracy respectively, under an imbalance factor of 100 using a single model.
-Finally AGLU, further boosts the performance of APA* by $1.1$pp and $0.4$pp for imbalance factor of 10 and 100 respectively.
+On iNaturalist18, Places-LT and CIFAR100-LT, our SE baselines with bag of tricks again reach the SOTA, as shown in Table III-a and b.
+Regarding, iNaturalist18, APA* improves the SE-baseline by 1.01.0pp.
+It also outperforms BCL [zhu2022balanced] by 0.50.5pp, ResLT by 1.81.8pp and LWS+ImbSAM [zhou2023imbsam] by 1.21.2pp.
+AGLU further enhances the performance by a staggering 3.53.5pp compared to SE-baseline, which is a significant increase.
+On Places-LT, our APA* also increases the performance of the SE baseline by 0.80.8pp.
+Moreover, it surpasses MisLAS [zhong2021improving] by 0.90.9pp and CC-SAM [zhou2023class] by 0.70.7pp.
+When AGLU is combined with APA*, it further increases the performance by 0.70.7pp reaching $42.0\%$.
+As shown in Table III-b, on CIFAR100-LT, APA* improves the performance by 0.50.5pp and 1.01.0pp compared to SE baseline using an imbalance factor of 10 and 100 respectively.
+Furthermore, compared to RIDE+CMO [park2022majority], CC-SAM [zhou2023class] and RIDE+CMO+CR [ma2023curvature], APA* achieves 1.91.9pp, 1.11.1pp and 1.21.2pp higher accuracy respectively, under an imbalance factor of 100 using a single model.
+Finally AGLU, further boosts the performance of APA* by 1.11.1pp and 0.40.4pp for imbalance factor of 10 and 100 respectively.
 Activation Ablation Study.
-We compare APA, without the Dropout and Layernorm, against different activation functions such as the Sigmoid and the Gumbel [1] using the SE-ResNet50 baseline.
+We compare APA, without the Dropout and Layernorm, against different activation functions such as the Sigmoid and the Gumbel [alexandridis2022long] using the SE-ResNet50 baseline.
 Also, we implement Sigmoid and Gumbel variants with learnable temperature, to further understand their difference to our adaptive activation.
-As shown in Table 4-a, Sigmoid with temperature achieves slightly better performance for the rare classes, however it over-fits the frequent categories.
-In contrast, our APA achieves the best performance, increasing the overall performance by $0.3$pp, the frequent classes by $0.2$pp and the common categories by $0.4$pp compared to the second best Gumbel with temperature.
+As shown in Table IV-a, Sigmoid with temperature achieves slightly better performance for the rare classes, however it over-fits the frequent categories.
+In contrast, our APA achieves the best performance, increasing the overall performance by 0.30.3pp, the frequent classes by 0.20.2pp and the common categories by 0.40.4pp compared to the second best Gumbel with temperature.
 Generalisation to other Attention mechanisms.
-In Table 4-b, we show that APA can be combined with other attention mechanisms.
-Specifically, APA improves the performance of Spatial Attention by $0.4$pp and the Spatial-Channel Attention by $1.3$pp.
-AGLU further increases the performance of Spatial Attention by $1.2$pp and Spatial-Channel Attention by $0.2$pp.
-Table 5: Comparisons on LVISv1.0 using MaskRCNN-FPN and 2x schedule.
+In Table IV-b, we show that APA can be combined with other attention mechanisms.
+Specifically, APA improves the performance of Spatial Attention by 0.40.4pp and the Spatial-Channel Attention by 1.31.3pp.
+AGLU further increases the performance of Spatial Attention by 1.21.2pp and Spatial-Channel Attention by 0.20.2pp.
+TABLE V: Comparisons on LVISv1.0 using MaskRCNN-FPN and 2x schedule.
 Method Backbone $AP^{m}$ $AP^{r}$ $AP^{c}$ $AP^{f}$ $AP^{b}$
-RFS  [22] R50 23.7 13.3 23.0 29.0 24.7
-IIF [2] 26.3 18.6 25.2 30.8 25.8
-Seesaw [89] 26.4 19.6 26.1 29.8 27.4
-LOCE [20] 26.6 18.5 26.2 30.7 27.4
-PCB+Seesaw [30] 27.2 19.0 27.1 30.9 28.1
-ECM  [39] 27.4 19.7 27.0 31.1 27.9
-GOL  [1] 27.7 21.4 27.7 30.4 27.5
-ECM+GAP [107] 26.9 20.1 26.8 30.0 27.2
+RFS  [gupta2019lvis] R50 23.7 13.3 23.0 29.0 24.7
+IIF [alexandridis2023inverse] 26.3 18.6 25.2 30.8 25.8
+Seesaw [wang2021seesaw] 26.4 19.6 26.1 29.8 27.4
+LOCE [feng2021exploring] 26.6 18.5 26.2 30.7 27.4
+PCB+Seesaw [he2022relieving] 27.2 19.0 27.1 30.9 28.1
+ECM  [hyun2022long] 27.4 19.7 27.0 31.1 27.9
+GOL  [alexandridis2022long] 27.7 21.4 27.7 30.4 27.5
+ECM+GAP [zhang2023reconciling] 26.9 20.1 26.8 30.0 27.2
 GOL (baseline) SE-R50 28.2 20.6 28.9 30.8 28.1
-GOL+AGLU(ours) APA*-R50 29.1${}^{\textbf{+0.9}}$ 21.6${}^{\textbf{+0.2}}$ 29.6${}^{\textbf{+0.7}}$ 31.7${}^{\textbf{+0.6}}$ 29.0${}^{\textbf{+0.9}}$
-RFS [22] R101 [28] 27.0 16.8 26.5 32.0 27.3
-NorCal [65] 27.3 20.8 26.5 31.0 28.1
-Seesaw [89] 28.1 20.0 28.0 31.8 28.9
-GOL [1] 29.0 22.8 29.0 31.7 29.2
-ECM [39] 28.7 21.9 27.9 32.3 29.4
-PCB + Seesaw[30] 28.8 22.6 28.3 32.0 29.9
-ROG [107] 28.8 21.1 29.1 31.8 28.8
+[…]
+RFS [gupta2019lvis] R101 [he2016deep] 27.0 16.8 26.5 32.0 27.3
+NorCal [pan2021model] 27.3 20.8 26.5 31.0 28.1
+Seesaw [wang2021seesaw] 28.1 20.0 28.0 31.8 28.9
+GOL [alexandridis2022long] 29.0 22.8 29.0 31.7 29.2
+ECM [hyun2022long] 28.7 21.9 27.9 32.3 29.4
+PCB + Seesaw[he2022relieving] 28.8 22.6 28.3 32.0 29.9
+ROG [zhang2023reconciling] 28.8 21.1 29.1 31.8 28.8
 GOL (baseline) SE-R101 29.7 23.0 29.9 32.5 30.0
-GOL+AGLU(ours) APA*-R101 30.7${}^{\textbf{+1.0}}$ 23.6${}^{\textbf{+0.6}}$ 31.3${}^{\textbf{+1.4}}$ 33.1${}^{\textbf{+0.7}}$ 31.11.${}^{\textbf{1.1}}$
+[…]
 Combining APA* with Classifier Learning.
-APA* is an efficient module that can be easily combined with common classifier learning techniques such as margin adjustment [71], reweighting [8] and resampling [44].
-As shown in Table 4-c, APA* consistently boosts the performance of all these methods.
+APA* is an efficient module that can be easily combined with common classifier learning techniques such as margin adjustment [Ren2020balms], reweighting [cao2019learning] and resampling [kang2019decoupling].
+As shown in Table IV-c, APA* consistently boosts the performance of all these methods.
 Larger ResNets.
 We further train deeper models like ResNet-101 and ResNet-152 on ImageNet-LT and compare the SE and APA* attention methods.
-As Table 4-d shows, the APA* module enlarges the performance of all models consistently for both frequent, common and rare classes.
-Especially, APA* with ResNet101 promotes overall accuracy by $2.1$pp, boosting the rare categories by a significant $4.5$pp compared to SE-Resnet101.
+As Table IV-d shows, the APA* module enlarges the performance of all models consistently for both frequent, common and rare classes.
+Especially, APA* with ResNet101 promotes overall accuracy by 2.12.1pp, boosting the rare categories by a significant 4.54.5pp compared to SE-Resnet101.
 Ablation of APA* components.
-APA* uses Channel attention, APA, Dropout and Layer-Norm and we show their effects in Table 4-e.
-The plain R50, without channel attention, achieves $55.0\%$ and SE increases its performance by $1.0$pp.
-APA increases the performance of the SE baseline by another $1.0$pp.
-Dropout increases the performance by $0.3$pp and LayerNorm enhances the performance by an additional $0.1$pp.
+APA* uses Channel attention, APA, Dropout and Layer-Norm and we show their effects in Table IV-e.
+The plain R50, without channel attention, achieves $55.0\%$ and SE increases its performance by 1.01.0pp.
+APA increases the performance of the SE baseline by another 1.01.0pp.
+Dropout increases the performance by 0.30.3pp and LayerNorm enhances the performance by an additional 0.10.1pp.
 In the Appendix, we show the full component ablation.
 Comparison of AGLU.
-In Table 4-f, we compare AGLU to other commonly used activations functions, using APA* ResNet50 as backbone and switching the activation function of the intermediate layers.
+In Table IV-f, we compare AGLU to other commonly used activations functions, using APA* ResNet50 as backbone and switching the activation function of the intermediate layers.
 Our AGLU outperforms all the other methods and it is the best choice.
-Long-Tailed Instance Segmentation.
-As the results suggest in Table 5, the SE-R50 backbone increases the overall mask performance compared to plain GOL-R50 by $0.5$pp but most improvement comes from the common and frequent categories while the rare categories are significantly reduced by $0.8$pp.
-In contrast, our APA* with AGLU-R50 improves the performance by $0.9$pp on average mask and bounding box precision, $1.0$pp on $AP_{r}$, $0.7$pp on $AP_{c}$ and $0.9$pp on $AP_{f}$ compared to SE-R50-GOL.
-Compared to GOL with SE-R101, APA* with AGLU also improves the performance by $1.0$pp on $AP$, $0.6$pp on $AP^{r}$, $1.4$pp on $AP^{c}$, $0.7$pp on $AP^{f}$ and $1.1$pp on $AP^{b}$.
+V-D Long-Tailed Instance Segmentation.
+As the results suggest in Table V, the SE-R50 backbone increases the overall mask performance compared to plain GOL-R50 by 0.50.5pp but most improvement comes from the common and frequent categories while the rare categories are significantly reduced by 0.80.8pp.
+In contrast, our APA* with AGLU-R50 improves the performance by 0.90.9pp on average mask and bounding box precision, 1.01.0pp on $AP_{r}$, 0.70.7pp on $AP_{c}$ and 0.90.9pp on $AP_{f}$ compared to SE-R50-GOL.
+Compared to GOL with SE-R101, APA* with AGLU also improves the performance by 1.01.0pp on $AP$, 0.60.6pp on $AP^{r}$, 1.41.4pp on $AP^{c}$, 0.70.7pp on $AP^{f}$ and 1.11.1pp on $AP^{b}$.
 This highlights that our APA* and AGLU modules are robust for the rare classes and they outperform the previous SOTA in long-tailed instance segmentation.
@@
 More visualisations are shown in Appendix.
-Qualitative Results.
+V-E Qualitative Results.
 We use APA* and Imagenet-LT for our qualitative analysis.
 In Fig.
-5-(a) we show that APA* increases the variance of channel attention for the most semantic layer by $0.04$ compared to the baseline, making it diverse and informative for all channels.
+6-(a) we show that APA* increases the variance of channel attention for the most semantic layer by $0.04$ compared to the baseline, making it diverse and informative for all channels.
 In Fig.
-5-(b), we show that APA* increases the entropy of the attention signal for the deeper layers by $0.4\sim 0.6$ compared to the baseline, showing that our module produces informative signals that effectively attend to the rare classes.
-In Fig.-c 5-top, we show that the baseline channel attention of the rare classes, produces all-pass filters in other words, attention signals that have small entropy, i.e., $E=0.018$ for the bee house and $E=0.098$ for the warthog respectively.
+6-(b), we show that APA* increases the entropy of the attention signal for the deeper layers by $0.4\sim 0.6$ compared to the baseline, showing that our module produces informative signals that effectively attend to the rare classes.
+In Fig.-c 6-top, we show that the baseline channel attention of the rare classes, produces all-pass filters in other words, attention signals that have small entropy, i.e., $E=0.018$ for the bee house and $E=0.098$ for the warthog respectively.
 This hinders rare class learning and it results in misclassification.
 In contrast, our method in Fig.
-5-bottom produces informative channel attention signals that have larger entropy i.e.
+6-bottom produces informative channel attention signals that have larger entropy i.e.
 $E=0.46$ and $E=0.91$ for the bee house and warthog classes respectively, allowing the model to retrieve rare class features and to make correct predictions.
 Visualisations of the learned $\kappa$ and $\lambda$ parameters are provided in the Appendix.
-Generalisation to Balanced tasks.
-In Table 6-a, we show that APA*+AGLU increases the performance of SE-MaskRCNN-R50 by $0.7$pp on $AP^{b}$ and $AP^{m}$ on COCO.
-In Table 6-b, we show that APA*+AGLU increase the performance of SE-FasterRCNN-R50 by $2.9$pp and the performance of SE-Cascade-RCNN-R50 by $2.1$pp on the challenging V3Det, which are significant increases considering that this dataset has 13K classes.
-In Table 6-c, we show that our methods increase the performance for R50, SE-R50 and CBAM-R50, by $0.6$pp, $1.2$pp and $0.6$pp respectively.
-Finally in Table  6-d, we show that our modules increase the performance by $1.2$pp on SE-R50, $0.9$pp on SE-R101 and $0.5$pp on SE-R101, showing that they can generalise in balanced training.
-Table 6: Resuls on balanced tasks.
-In (a) and (b), we combine APA*+AGLU on COCO dataset and V3Det dataset respectively.
-In (c) we combine AGLU and APA* to ResNet50, SE-ResNet50 and CBAM-ResNet50 in ImageNet1K.
-In (d) we show that APA*+AGLU generalises to deeper backbones in ImageNet1K.
+VI Results on other tasks
+In this section, we comprehensively evaluate the generalisation ability of AGLU activation beyond long-tail learning.
+We test AGLU in balanced classification, object detection, large-scale supervised pretraining and finetuning, Multi-modal LLMs, image generation with diffusion transformer and pure language models.
+In every subsection, we introduce the model architecture, training details, dataset and discuss the results.
+VI-A Image Classification
+We perform experiments on ImageNet1K [deng2009imagenet], using ResNets [he2016deep] and ViT [dosovitskiy2021an] models.
+We train Resnets with channel attention [hu2018squeeze], and spatial attention [woo2018cbam] for 100 epochs using learning rate 0.256, batch size 256, weight decay 6e-6, mixup coefficient 0.2, AutoAugment [cubuk2019autoaugment], training crop size 224 and random-resized-crops.
+We adopt ViT-S and B variants using a convolutional patch embedding [xiao2021early], layer-scale [touvron2021going], cosine positional embeddings and average pooling instead of class token [beyer2022better].
+We train these models for 300 epochs using learning rate 0.002, weight decay 0.05, mixup coefficient 0.2, cutmix coefficient 1.0, batch size 2048, RandAugment [cubuk2020randaugment], repeated augmentations, color jittering, training crop size 192, random-resized-crops.
+For ViT-B, we add stochastic depth [huang2016deep] 0.2 to prevent overfitting.
+As Table VI-a shows, our method increases the accuracy for R50, SE-R50 and CBAM-R50, by 0.60.6pp, 1.21.2pp and 0.60.6pp respectively.
+Also, it increases accuracy by 0.90.9pp on SE-R101 and 0.50.5pp on SE-R152 showing that AGLU is robust for any model size.
+In Table VI-b we show that AGLU has competitive performance for ViT models, increasing by 0.20.2pp and 0.30.3pp the ViT-S and ViT-B models respectively.
+TABLE VI: Results on ImageNet1K using ResNets in (a) and vision Transformers in (b).
+Method top-1
+ResNet50 [hu2018squeeze] 76.9
+w/ AGLU 77.5
+SE-ResNet50 [hu2018squeeze] 77.5
+w/ APA*+AGLU 78.7
+CBAM-ResNet50 [hu2018squeeze] 78.3
+w/ APA*+AGLU 78.9
+SE-ResNet101 79.4
+w/ APA* + AGLU 80.3
+SE-ResNet152 80.3
+w/ APA* + AGLU 80.8
+Method top-1
+ViT-S 80.2
+w/ AGLU 80.4
+ViT-B 82.5
+w/ AGLU 82.8
+(a) Using ResNets
+(b) Using ViT
+TABLE VII: Dataset details of downstream tasks.
+Dataset Train size Test size Number of Classes
+ImageNet[deng2009imagenet] 1,281,167 50,000 1000
+iNaturalist2017 [van2018inaturalist] 579,184 95,986 5,089
+iNaturalist2018 [van2018inaturalist] 437,513 24,426 8,142
+iNaturalist2021 [van2018inaturalist] 2,686,843 500,000 10,000
+CIFAR10 [krizhevsky2009learning] 50,000 10,000 10
+CIFAR100 [krizhevsky2009learning] 50,000 10,000 100
+Places365 [zhou2014learning] 1,803,459 36,499 365
+Training on ImageNet21K.
+We test AGLU under large scale training by performing ImageNet21k pretraining and finetuning.
+For this experiment we use a SE-ResNet50 and a simple ViT-B architecture that strictly follows [beyer2022better].
+We train these models using batch size 2048, learning rate 0.004, Adam with momentum [kingma2014adam], weight decay 0.02, mixup coefficient 0.2, training crop size 224, RandAugment and random-resized-crop for 90 epochs.
+After pretraining, we perform finetuning on 7 downstream tasks, including finegrained classification, scene classification and object classification, using the datasets in Table VII.
+In more detail, we finetune the models for 50 epochs with learning rate 0.04, SGD, no weight decay, batch size 512, mixup coefficient 0.2, cutmix coefficient 1.0, RandAugment, color-jittering, label-smoothing 0.1, train-crop 224 and random-resized-crops.
+TABLE VIII: Comparative downstream results on 7 popular image classification datasets using 224 training and validation crops.
+Method Imagenet1K iNaturalist 2017 iNaturalist 2018 iNaturalist 2021 CIFAR10 CIFAR100 Places365 Average
+SE-Resnet50 80.5 67.3 73.4 82.1 96.0 86.4 54.3 77.1
+w/ APA* + AGLU 81.5 69.7 76.4 83.9 97.6 87.6 54.7 78.8
+ViT-B 83.9 72.1 78.9 84.7 99.1 93.5 56.9 81.3
+w/ AGLU 84.0 72.2 79.3 85.1 99.2 93.6 57.1 81.5
+The results are shown in Table VIII.
+Our activation significantly enhances the performance of SE-ResNet50 by 1.91.9pp macro-average.
+AGLU, also increases the performance of VIT-B by 0.20.2pp macro average, showing that AGLU is a good activation function for both Transformer and CNN architectures.
+VI-B Experiments with Object detection.
+We perform experiments on COCO [lin2014microsoft], Objects365v2[shao2019objects365] and the recently proposed V3Det[wang2023v3det], which is a challenging large scale detection dataset with 13K classes and 243K images.
+We report $AP^{b}$ and $AP^{m}$ for COCO and $AP^{b}$ for V3Det and Objects365v2, using primarily SE-ResNets backbones pretrained on ImageNet1K.
+We use the mmdetection framework and train all models using 2x schedule.
+In Table IX-a, we show that APA*+AGLU increases the COCO performance of SE-MaskRCNN-R50 by 0.70.7pp on $AP^{b}$ and $AP^{m}$.
+Also, it increases the performance of SE-MaskRCNN-R101 by 0.90.9pp on $AP^{b}$ and 0.50.5pp on $AP^{m}$.
+As Table IX-b shows, APA*+AGLU improves $AP^{b}$ by 1.11.1pp and 1.61.6pp on the vast Objects365v2 dataset using the SE-FasterRCNN-R50 and R101 models respectively.
+In Table IX-c, we show that APA*+AGLU increase the performance of SE-FasterRCNN-R50 by 2.92.9pp and the performance of SE-Cascade-RCNN-R50 by 2.12.1pp on the challenging V3Det, which are significant increases considering that this dataset has 13K classes.
+TABLE IX: Resuls on balanced tasks.
+In (a) and (b) and (c), we combine APA*+AGLU on $CO$, Objects365v2 and V3Det dataset respectively.
+In (d) we combine AGLU and APA* to ResNet, SE-ResNet and CBAM-ResNet in ImageNet1K.
+The shows that the baseline is taken from mmdetection [mmdetection].
 Method $AP^{b}$ $AP^{m}$
-MaskRCNN-R50 39.2 35.4
-w/ SE 40.5 36.9
+MaskRCNN-ResNet50 39.2 35.4
+SE-MaskRCNN-ResNet50 40.5 36.9
 w/ APA*+AGLU 41.2 37.6
+SE-MaskRCNN-ResNet101 42.7 38.4
+w/ APA*+AGLU 43.6 38.9
 Method $AP^{b}$
-FasterRCNN-R50 25.4
+FasterRCNN-ResNet50 19.8
+SE-FasterRCNN-ResNet50 26.4
+w/ APA*+AGLU 27.5
+SE-FasterRCNN-ResNet101 29.3
+w/ APA*+AGLU 30.9
+Method $AP^{b}$
+FasterRCNN-Resnet50 25.4
 w/ SE 27.0
 w/ APA*+AGLU 29.9
-CascadeRCNN-R50 31.6
+CascadeRCNN-ResNet50 31.6
 w/ SE 33.3
-APA*+AGLU-CRCNN 35.4
-Method top-1
-ResNet50 [37] 76.9
-w/ AGLU 77.5
-SE-ResNet50 [37] 77.5
-w/ APA*+AGLU 78.7
-CBAM-ResNet50 [37] 78.3
-w/APA*+AGLU 78.9
-Method top-1
-SE-R50 77.5
-+APA* 77.9
-+APA* + AGLU 78.7
-SE-R101 79.4
-+APA* 79.2
-+APA* + AGLU 80.3
-SE-R152 80.3
-+APA* 80.5
-+APA* + AGLU 80.8
+w/ APA*+AGLU 35.4
 (a) COCO
-(b)V3Det
-(c)ImageNet1k
-(d) ImageNet1k
-6 Related Work
-Long-tailed image recognition.
-Long-tailed image recognition can be grouped according to representation learning and classifier learning techniques.
-Representation learning techniques improve the feature quality through rare class features generation [85, 87, 104], contrastive objectives [91, 15, 117, 74, 53, 43], ensemble or fusion models [95, 51, 14, 113, 50, 13, 7], knowledge distillation [51, 41, 29, 52], knowledge transfer [58, 67, 118] and data augmentation [111, 68, 101].
-Classifier learning techniques enhance rare class classification through decoupled training [44, 108, 92, 47, 36], margin adjustment [63, 71, 34, 8, 39, 65, 89, 93, 110, 2], cost-sensitive learning [16, 46, 96, 38], resampling [9, 61, 76, 120, 22, 68, 33, 104], dynamic loss adaptation based on batch statistics [35, 80, 94], gradient statistics [79, 49], weight norms [93] and classification scores [20, 30].
-These techniques are efficient, however they are applied either during the input phase or during the loss disregarding the intermediate channel activations.
-Attention Networks.
-Attention networks, such as spatial attention [42], channel attention [37] and spatial-channel attention [98], are widely used in image recognition.
-Self-attention is the core mechanism of the Visual Transformer (ViT) architecture [19], and recently many improvements have been made to the Transformer architecture [81, 83, 56] and its training procedure [78, 6, 82].
-In our work, we primarily used APA with channel attention models, but it also works with plain ResNets, Spatial and Spatial-Channel attention models.
-In the Appendix, we show results with self-attention models.
-Activation functions The RELU function [21] dominates the landscape of deep image classification, and it is especially used inside the convolutional networks, while the GELU and SiLU [31] are more commonly used inside the transformer network [19] or the ConNext models [57, 97].
-The PRELU function [27] is a generalisation of the RELU, because it linearly activates both positive and negative inputs and the ELU [11] is a follow-up work that enforces saturation on the negative inputs after some threshold.
-Our AGLU activation, generalises the RELU, the GELU and SiLU and it can also be seen as a smoother version of PRELU.
-7 Conclusion
+(b) Objects365v2
+(c) V3Det
+VI-C Experiments with Vision Language Model.
+We further test the performance of our proposed activation function in the visual instruction finetuning task using vision-language models (VLMs), following the LLAVA series paradigm [li2024llava, liu2023llava, liu2024llavanext, li2024llavanext-ablations, li2024llavanext-strong].
+In more detail, we use an MLP with AGLU activation as a connector module and pretrain two VLMs, including the Qwen2 [yang2024qwen2technicalreport] and Deepstack [meng2024deepstack] with the SigLIP-so400M [zhai2023sigmoid] vision encoder, using the LLaVA-1.5-PT dataset that has 558k image-text pairs [liu2023improvedllava], keeping all parameters frozen except for the connector MLP.
+Next, we perform instruction tuning and we fully finetune the models using the OpenLLaVA-Next dataset [chen2024open].
+To evaluate the performance, we use the LMM-evaluation kit [zhang2024lmmsevalrealitycheckevaluation] that includes many vision-language benchmarks such as MME [fu2023mme], AI2D [kembhavi2016diagram], DocVQA [mathew2021docvqa], GQA [hudson2019gqa], MMB [liu2024mmbench], OCRB [Liu_2024_ocrbench], MMMU [yue2024mmmu], OKVQA [marino2019ok], POPE [li2023evaluating], SQA [lu2022learn], TextVQA [singh2019towards] and VisWiz [gurari2018vizwiz].
+As shown in Table X, our AGLU connector increases the performance, in both Qwen2-1.5B, Deepstack-7B and Qwen2-7B models by 2.42.4pp, 0.60.6pp and 0.90.9pp, respectively, on the macro average metric and achieves superior performance compared to open source models like LLaVA-Next-7B [liu2024llavanext] and proprietary models like DeepSeek-Janus-Pro 7B.
+This indicates that AGLU is particularly useful in vision language training because its adaptive behaviour can bridge the domain gap between language and image distributions.
+TABLE X: Comparative results on the visual instruction following task, using a Qwen2-SigLIP model.
+[…]
+DeepSeek-Janus-Pro-7B [chen2025janus] 1567 - - 59.1 - - 79.2 - 41.0 - - 87.4 - - - -
+LLaVA-1.5-7B [liu2023improvedllava] 1511 - - 62.0 - 58.3 64.3 - - - - 86.0 71.6 61.3 50.0 -
+LLaVA-Next-7B [liu2024llavanext] 1519 66.6 74.4 64.2 - 60.6 67.4 - 35.8 - - 86.5 70.1 64.9 57.6 -
+VILA-7B [lin2024vila] 1533 - - 62.3 - 61.7 70.3 - - - - 85.5 68.3 64.4 57.8 -
+InternVL-Chat-7B [chen2024internvl] 1525 - - 62.9 - - - - - - - 87.1 - 57.0 52.5 -
+Qwen2-1.5B-SigLIP 1393 59.3 46.6 60.3 33.7 57.3 57.6 33.8 27.9 42.5 51.4 86.5 70.3 53.4 37.5 52.5
+w/ AGLU 1390 60.3 47.7 60.9 37.7 60.1 64.3 47.6 26 41.3 52.3 87.5 69.8 54.9 43.4 54.9
+Deepstack-7B [meng2024deepstack] 1592 74.6 73.2 64.4 46.7 72.8 75.3 51.0 47.1 53.7 56.2 88.0 71.4 66.1 61.7 65.5
+w/ AGLU 1573 75.3 73.6 64.3 47.3 73.6 76.3 51.3 49.0 53.0 58.0 87.9 76.1 66.4 61.0 66.1
+Qwen2-7B-SigLIP 1596 73.9 74.8 63.9 42.9 72.3 74.2 56.3 45.8 55.9 61.2 87.7 70.6 68.0 53.8 65.4
+w/ AGLU 1634 74.1 75.6 64.3 46.9 71.8 75.1 55.1 45.7 58.3 61.9 88.1 76.5 68.0 51.0 66.3
+VI-D Experiments in Image Generation.
+We try our AGLU activation in the task of image generation.
+Specifically, we select LightningDIT-XL [yao2025reconstruction], which is a strong transformer-based, latent diffusion model as our baseline, and we replace all its activation functions with AGLU, keeping any other training details and hyperparameters fixed.
+We train the AGLU-LightiningDIT-XL model with patch size one, on the Imagenet1k-256 generation benchmark for 64 epochs and evaluate the generation quality using the ADM evaluation suite [dhariwal2021diffusion].
+As Table XI shows our activation function achieves good performance, decreasing the FID score by $0.1$ and increasing the Recall by 0.80.8pp compared to LightningDiT baseline.
+TABLE XI: AGLU achieves good performance in image generation.
+Model FID sFID Inception Score Precision Recall
+LightiningDIT-XL[yao2025reconstruction] 2.2 4.3 251.9 81.5 58.3
+LightiningDIT-XL w/ AGLU 2.1 4.2 251.1 81.0 59.1
+VI-E Experiments with Language Model.
+We perform a language next-token prediction experiment using GPT2 [radford2019language] and the FineWeb-Edu [lozhkov2024fineweb-edu] subset that contains 10 billion GPT2 tokens.
+The model implementation follows nanogpt by [Karpathy2024].
+We train four GPT2 models, GPT2-small that has 124 million parameters, GPT2-medium that has 350 million, GPT-large that has 774 million parameters and GPT-xl that has 1.5 billion parameters for 10 billion tokens, with a batch size of 0.5 million tokens, learning rate 6e46e-4 and Adam optimizer with momentum [kingma2014adam].
+We test the model on the HellaSwag benchmark [zellers2019hellaswag] using zero-shot evaluation.
+To apply AGLU with GPT2, we simply replace all GELU activations with AGLU inside all MLP layers of the transformer.
+As the results suggest in Table XII, AGLU increases the performance of GPT2 by 0.2pp, 0.4pp, 0.3pp and 0.8pp on the HellaSwag benchmark using GPT2-small, medium, large and xl respectively, and reduces the validation loss on the FineWeb-Edu by 0.01 and 0.02 on GPT2-medium, and GPT2-large and xl respectively showing that AGLU is a good alternative for the next token prediction task, especially for deeper models.
+TABLE XII: Comparative results using GPT2 model on the HellaSwag benchmark and Fineweb-Edu validation set.
+Method HellaSwag FineWeb Val-Loss
+GPT2-small 30.8 3.08
+GPT2-small with AGLU 31.0 3.08
+GPT2-medium 35.2 2.86
+GPT2-medium with AGLU 35.6 2.85
+GPT2-large 38.8 2.72
+GPT2-large with AGLU 39.1 2.70
+GPT2-xl 42.7 2.62
+GPT2-xl with AGLU 43.5 2.60
+VII Analysis
+VII-A Computation cost.
+AGLU costs more computing resources compared to non-parametric activation functions because it requires training.
+The additional computation depends on the architecture, the optimizer and the size of the layer where AGLU is applied.
+In Table XIII, we compare AGLU against the vanilla architecture for different models, in terms of the number of parameters, peak memory measured in MB, throughput measured in samples per second and FLOPS using ImageNet1K.
+AGLU requires $17\%$ more memory, it has $3\%$ smaller throughput and it has the same number of parameters and FLOPS in the SE-ResNet50 architecture.
+In the ViT-B architecture, it lowers the throughput by $3\%$ but keeps all other measures the same.
+The extra computing occurs mostly during training and it depends on the network’s architecture, i.e., if AGLU is applied after a small linear layer, then the training cost is low, however, if it is applied after a large layer, i.e.
+a large convolution layer, then the training cost is larger.
+VII-B Learned activations.
+We visualise the converged AGLU activation functions, i.e.
+the output $AGLU(z,\lambda,\kappa)=y$ for the ImageNet21K-ViT-B and GPT models.
+To better understand their behaviour, we measure AGLU’s expected output $E[y]$, for random inputs $z$ drawn from the Normal distribution.
+It is striking to notice that both the ViT and GPT2, have the same activation patterns, i.e.
+the activation of the early layers is harder, i.e., more non-linear and it allows mostly positive inputs to pass whereas in the deeper layers it becomes smoother, i.e.
+more linear, allowing both positive and negative outputs to propagate.
+We hypothesize that, given the models’ hierarchical feature extraction ability, the network distinguishes the early features easily and draw harder decision boundaries with the activation function, but as the layers grow deeper, the features become more abstract and the network distinguishes the features less easily and uses smoother decision boundaries.
+Figure 7: Comparison of the expected AGLU-outputs $E[y]$ for ViT and GPT models, given input drawn from the Normal distribution, across layers.
+The expected value is larger at the first layers, indicating harder activations, and it becomes smaller at the last layers, indicating smoother activations, for both GPT and ViT.
+TABLE XIII: Computation cost of AGLU.
+FLOPS are measured on a V100 GPU using batch-size of 1, $224^{2}$ image resolution and mixed precision.
+Peak memory (MB) and throughput ($\frac{\text{img}}{\text{sec}}$) are measured on the ImageNet1K finetuning.
+Model Parameters MB $\frac{\text{img}}{\text{sec}}$ GFLOPS
+SE-R50 28M 6020 393.6 8.2
+w/ APA*+AGLU 28M 7064 382.7 8.2
+ViT-B 86M 5145 263.7 33.5
+w/ AGLU 86M 5145 256.9 33.5
+VIII Discussion
+We have shown that the adaptive activation function is a solid alternative to standard activation functions such as RELU or GELU, for many vision tasks, especially for long-tailed tasks.
+Some of the biggest benefits are that, it is a generalisation of many common activation functions, it can be computed analytically and it is a easy to use.
+We believe that AGLU, can be particularly effective when applied to specialised modules such as in the VLM connector, or CNN detectors and is a good alternative for large scale pretrained Transformer models, showing 0.30.3pp increase in ViT and GPT models for both classification and generation tasks.
+This is lower than the performance increase observed with CNNs, potentially because the transformer is a highly non-linear structure and already very expressive due to its attention mechanism.
+Regarding, the computational resources, AGLU is overall more expensive than non-adaptive functions but the computing cost is diminishing when AGLU is applied to small linear layers.
+IX Conclusion
 Our work highlights the impact of the activation function inside the model’s activations for balanced and imbalanced data distributions.
-We have empirically showed that the degree of data imbalance affects the logit distributions and the intermediate signals and we have showed that the commonly used Sigmoid activation function is unable to model the intermediate features.
-To this end, we have proposed an novel adaptive parametric activation that unifies most common activation functions under the same formula and we have tested it in several long-tail and balanced classification and detection benchmarks showing great generalisation.
-Acknowledgements. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) project “ViTac: Visual-Tactile Synergy for Handling Flexible Materials” (EP/T033517/2).
-Appendix 0.A Representations’ Quality
-We evaluate the quality of the representations learned by the SE and APA* models using the recently proposed Neural Collapse framework [66].
+We have empirically shown that the degree of data imbalance affects the logit distributions and the intermediate signals and we have shown that the commonly used Sigmoid activation function is unable to model the intermediate features.
+To this end, we have proposed an novel adaptive parametric activation that unifies most common activation functions under the same formula and we have tested it in several long-tail and balanced benchmarks showing great generalisation.
+Appendix A Representations’ Quality
+We evaluate the quality of the representations learned by the SE and APA* models using the recently proposed Neural Collapse framework [papyan2020prevalence].
 Let $f_{k,j}\in\mathrm{R^{d}}$ be the features of the penultimate layer, $k=\{1,2,...,K\}$ the class, $n_{k}$ the number of samples in the class $k$ and $n=\sum_{k=1}^{K}n_{k}$ the total number of samples in the dataset.
@@
 The within-class covariance matrix $\Sigma_{W}\in\mathrm{R^{d\times d}}$ and between-class covariance matrix $\Sigma_{B}\in\mathrm{R^{d\times d}}$ are:
-$\begin{split}\Sigma_{W}:=&\frac{1}{n}\sum_{k=1}^{K}\sum_{j=1}^{n_{k}}(f_{k,j}-\bar{f_{k}})(f_{k,j}-\bar{f_{k}})^{\top}\\ \Sigma_{b}:=&\frac{1}{K}\sum_{k=1}^{K}(\bar{f_{k}}-f_{G})(\bar{f_{k}}-f_{G})^{\top}\\ \end{split}$ (11)
+[…]
 The $\Sigma_{W}$ matrix shows how distant are individual features $f_{k,j}$ from their class prototype $\bar{f_{k}}$ and it is an indicator of feature compactness.
 The $\Sigma_{b}$ matrix shows how distant are the class prototypes from the global feature, indicating the class separability.
-Using these matrices we measure the Neural collapse Variability $NC1$ according to [119] as follows:
+Using these matrices we measure the Neural collapse Variability $NC1$ according to [zhu2021geometric] as follows:
 $NC1:=\frac{1}{K}trace(\Sigma_{W}\Sigma_{b}^{\dagger})$ (12)
 where the $\dagger$ symbol denotes the pseudo inverse of $\Sigma_{b}$.
-$NC1$ measures the magnitude of the within-class covariance $\Sigma_{W}$ compared to the magnitude of the between-class covariance $\Sigma_{b}$ as explained in [119].
+$NC1$ measures the magnitude of the within-class covariance $\Sigma_{W}$ compared to the magnitude of the between-class covariance $\Sigma_{b}$ as explained in [zhu2021geometric].
 In practise, a low $NC1$ measure shows that the model has more compact features since $\Sigma_{W}\downarrow$ decreases and more separable class prototypes because the $\Sigma_{b}\uparrow$ increases.
-Having more compact features and more separable class prototypes make the representations better and enhance the classification results as shown empirically in previous works [112, 102, 103, 99].
-Using Equation 12, we measure the $NC1$ of the deep features of the penultimate layer of SE and APA*, in Table 7 using ImageNet-LT test-set.
+Having more compact features and more separable class prototypes make the representations better and enhance the classification results as shown empirically in previous works [Zhong_2023_CVPR, yang2022inducing, yang2023neural, xie2023neural].
+Using Equation 12, we measure the $NC1$ of the deep features of the penultimate layer of SE and APA*, in Table XIV using ImageNet-LT test-set.
 As the results suggest, our APA* has lower $NC1$ measure for all backbones, showing that APA* produces superior representations that are more compact and seperable than the baseline.
 This provides another qualitative explanation why our APA* has better performance than SE.
-Table 7: Neural Collapse NC1 measure, on ImageNet-LT test set. APA* has lower $NC1$ measure than the baseline, which indicates that it has learned superior representations.
+TABLE XIV: Neural Collapse NC1 measure, on ImageNet-LT test set. APA* has lower $NC1$ measure than the baseline, which indicates that it has learned superior representations.
 Backbone SE-$NC1\downarrow$ APA*-$NC1\downarrow$
@@
 ResNet-152 3.24 2.69
-Appendix 0.B Implementation Details
-The implementation details of APA* and AGLU are shown in Table 8.
-For balanced ImageNet-1K, the $\lambda$ parameters are initialised as random variables drawn from a Uniform distribution (U), with low parameter 00, and high parameter $1.0$.
-The APA $\kappa$ parameters are initialised with $U(0,1)$ and the AGLU $\kappa$ parameters are with initialised with $U(0.8,1.2)$.
+Appendix B Implementation Details
+The implementation details of APA* and AGLU are shown in Table XV.
+For balanced ImageNet-1K, the $\lambda$ parameters are initialised as random variables drawn from a Uniform distribution (U), with low parameter 0, and high parameter $1.0$.
+The APA $\kappa$ parameters are initialised with $U(0,1)$ and the AGLU $\kappa$ parameters are with initialised with $U(1,1.3)$.
 For all other downstream tasks, that use a pretrained model, such as COCO, LVIS, Places-LT and V3Det, we don’t re-initialise the $\kappa$ and $\lambda$ parameters and we simply load them from the pretained ImageNet1K model.
-0.B.1 Stable APA implementation
+We do not use weight decay for the $\kappa$ and $\lambda$ parameters as in [he2015delving].
+B-A Stable APA implementation
 During the development of APA, we found that it is more stable to use Softplus $s_{f}(z,\beta)=\frac{1}{\beta}\ln(1+\exp(\beta z))$, than double exponents, when computing APA.
@@
 and it is equivalent to the APA used in the main paper.
+TABLE XV: Implementation details for Long-tailed Datasets, across various architectures.
 Method ImageNet-LT i-Naturalist18 Places-LT C100-LT LVISv1
@@
 Mixup $\alpha$ 0.2 0.2 0.2 0.2 -
-CutMix $\alpha$ 1.0 1.0 - -
-Label smoothing $\epsilon$ 0.1 0.1 - -
+CutMix $\alpha$ - 1.0 1.0 - -
+Label smoothing $\epsilon$ - 0.1 0.1 - -
 Repeated Aug - - - -
@@
 APA $\lambda$ Init U(0,1) U(0,1) N/A U(0,1) N/A
-AGLU $\kappa$ Init U(0.8,1.2) U(0.8,1.2) N/A U(0.8,1.2) N/A
+AGLU $\kappa$ Init U(1,1.3) U(1,1.3) N/A U(1,1.3) N/A
 AGLU $\lambda$ Init U(0,1) U(0,1) N/A U(0,1) N/A
-Table 8: Implementation details for Long-tailed Datasets, across various architectures.
-0.B.2 AGLU derivatives
+B-B AGLU derivatives
 Proof of Eq. 9. Then the gradient of AGLU with respect to $\kappa$ is:
@@
 $=x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\lambda x\exp(-\kappa x))$
-$=x^{2}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{-1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
+$=x^{2}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
 $=x^{2}\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
@@
 Proof of Eq. 11. Then the gradient of AGLU with respect to $\lambda$ is:
-$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial x}$ $=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$ (16)
-$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\cdot\partial\frac{(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$
-$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\kappa\lambda\exp(-\kappa x))$
-$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
-$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
-Appendix 0.C Results
-0.C.1 Experiments with AGLU and plain ResNets
-In Table 9, we show the result of AGLU when it applied inside plain ResNet50 models, (i.e.
-without channel attention).
-As the Table suggests, by simply replancing the RELU with AGLU, our method consistently increases the performance of plain ResNet models.
-Table 9: Top-1 accuracy on long-tailed classification datasets using ResNets.
-Dataset CIFAR100-LT ImageNet-LT iNaturalist
-Imbalance factor 10 100 256 500
-Model ResNet-32 ResNet50 ResNeXt50 ResNet50
-RELU 65.7 51.8 55.0 57.0 69.9
-AGLU (ours) 66.8 52.0 56.0 57.6 72.4
-0.C.2 Experiments with AGLU and Vision Transformers on ImageNet1K
-We perform a preliminary experiment with Vision Transformer models such as ViT [19] and Swin [56] using ImageNet1K.
-We replace the GELU activation with AGLU in every feedforward layer and we keep all other settings the same.
-As shown in Table 10, AGLU performs comparably to GELU.
-We believe this is because the Self-Attention function makes the features smooth, by removing their harmonising components and it makes them more Gaussian-like [90, 114, 18, 5].
-Consequently, the Gaussial linear error unit, GELU, might be a good choice for the ViT network and our AGLU method has comparable performance.
-Table 10: Results of AGLU using ViT models on ImageNet1K.
-Model Activation epochs top-1
-ViT-B [19] GELU 200 78.3
-ViT-B [19] AGLU 200 78.5
-Swin-T [56] GELU 100 78.7
-Swin-T [56] AGLU 100 78.9
-0.C.3 Impact of Initialisation
-In all of our experiments, we have initialised $\lambda$ using the Uniform distribution with low parameter 0 and high parameter 1 as a default.
-Regarding the $\kappa$ parameter inside AGLU, we initialise it to be close to $1.0$, as this works best, as shown in Table 11.
-Regarding the $\kappa$ parameter inside the attention layer, we found that initialising it with $Uniform(-1,0)$ is slightly better than $Uniform(0,1.0)$ as shown in Table 12.
-Table 11: AGLU-$\kappa$ parameter initialisation, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
+$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial x}=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+x\cdot\partial\frac{(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\kappa\lambda\exp(-\kappa x))\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}\\$ (16)
+Appendix C Results
+C-A Impact of Initialisation
+In all long-tail experiments, we have initialised $\lambda$ using the Uniform distribution with low parameter 0 and high parameter 1 as a default.
+Regarding the $\kappa$ parameter inside AGLU, we initialise it to be close to $1.0$, as this works best, as shown in Table XVI.
+Regarding the $\kappa$ parameter inside the attention layer, we found that initialising it with Uniform(1,0)Uniform(-1,0) is slightly better than $Uniform(0,1.0)$ as shown in Table XVII.
+In all other experiments including image generation, ImageNet21K, ImageNet1K training, VLM training and GPT training we used $\kappa$ and $\lambda$ initialization from the $Uniform(0.8,1.2)$ distribution, to simplify the codebase.
+The latter initialization has good performance across many balanced tasks.
+We note that it might be possible to achieve better performance by tuning the initialisation for every task, however we chose not to do it due to computing constrains.
+TABLE XVI: AGLU-$\kappa$ parameter initialisation, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
 AGLU - $\kappa$ top-1
 $Uniform(0,1)$ 57.7
-$Uniform(-2,0)$ 57.3
-$Uniform(-3,0)$ 57.4
-$Uniform(-2,2)$ Failed
-$Uniform(0.8,1.2)$ 57.9
-Table 12: $\kappa$ parameter initialisation inside the attention layer, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
+Uniform(2,0)Uniform(-2,0) 57.3
+Uniform(3,0)Uniform(-3,0) 57.4
+Uniform(2,2)Uniform(-2,2) Failed
+$Uniform(1,1.3)$ 57.9
+TABLE XVII: $\kappa$ parameter initialisation inside the attention layer, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
 APA - $\kappa$ top-1
 $Uniform(0,1)$ 57.6
-$Uniform(-1,0)$ 57.9
-0.C.4 Channel specific $\lambda$ and $\kappa$
+Uniform(1,0)Uniform(-1,0) 57.9
+C-B Channel specific $\lambda$ and $\kappa$
 We have also tried a variant that uses seperate $\lambda$ and $\kappa$ parameters for every channel.
-As shown in Table 13, this variant performs worse than using shared $\lambda$ and $\kappa$ parameters for the channels.
-Table 13: Results with Channel Specific $\lambda$ and $\kappa$, using APA* ResNet50 backbone on ImageNet-LT.
+As shown in Table XVIII, this variant performs worse than using shared $\lambda$ and $\kappa$ parameters for the channels.
+TABLE XVIII: Results with Channel Specific $\lambda$ and $\kappa$, using APA* ResNet50 backbone on ImageNet-LT.
 APA top-1
@@
 Channel Shared 57.9
-0.C.5 Baseline enhancements
-We show the detailed ablation study for ImageNet-LT in Table 14.
+C-C Baseline enhancements
+We show the detailed ablation study for ImageNet-LT in Table XIX.
 First, the vanilla ResNet50 model trained for 100 epochs on ImageNet-LT achieves $44.4\%$.
-When we train for 200 epochs then it adds $1.5$pp and switching from linear classifier to cosine classifier adds another $0.4$pp.
-Stronger training techniques such Mixup [106], Auto-Augment [12] and weight decay tuning further boost the performance by $3.3$pp.
-Post-calibrated Softmax [34] adds an additional $5.4$pp and finally the Squeeze and Excite module [37] adds another $1.0$pp reaching the final $56.0\%$.
+When we train for 200 epochs then it adds 1.51.5pp and switching from linear classifier to cosine classifier adds another 0.40.4pp.
+Stronger training techniques such Mixup [zhang2017mixup], Auto-Augment [cubuk2019autoaugment] and weight decay tuning further boost the performance by 3.33.3pp.
+Post-calibrated Softmax [hong2021disentangling] adds an additional 5.45.4pp and finally the Squeeze and Excite module [hu2018squeeze] adds another 1.01.0pp reaching the final $56.0\%$.
 Most baseline performance comes from the PC-Softmax and the weight decay finetuning.
-On top of this strong baseline, our APA increases the performance by $1.0$pp, showing its strong generalisability.
-Dropout and LayerNorm further increase the performance by $0.4$pp and finally AGLU adds a respectable $0.5$pp reaching $57.9\%$ accuracy on ImageNet-LT.
-The absolute improvement of all modules is $13.5$pp and our proposed methods, APA and AGLU, contribute by $1.5$pp which is a relative $11\%$ of the total absolute improvement.
-Table 14: Detailed Ablation Study, using ResNet50 on ImageNet-LT.
-200 epochs Cosine Classifier SE-nets[37] AutoAugment[12] Mixup [106] Weight Decay Tuning [3] PCS [34] APA (ours) Dropout [32] LayerNorm [4] AGLU (ours) ImagetNet-LT
+On top of this strong baseline, our APA increases the performance by 1.01.0pp, showing its strong generalisability.
+Dropout and LayerNorm further increase the performance by 0.40.4pp and finally AGLU adds a respectable 0.50.5pp reaching $57.9\%$ accuracy on ImageNet-LT.
+The absolute improvement of all modules is 13.513.5pp and our proposed methods, APA and AGLU, contribute by 1.51.5pp which is a relative $11\%$ of the total absolute improvement.
+TABLE XIX: Detailed Ablation Study, using ResNet50 on ImageNet-LT.
+200 epochs Cosine Classifier SE-nets[hu2018squeeze] AutoAugment[cubuk2019autoaugment] Mixup [zhang2017mixup] Weight Decay Tuning [alshammari2022long] PCS [hong2021disentangling] APA (ours) Dropout [hinton2012improving] LayerNorm [ba2016layer] AGLU (ours) ImagetNet-LT
 44.4
@@
 57.9
-0.C.6 Qualitative Results
+C-D Qualitative Results
 In Figure 8, we show the learned parameters, when training with the balanced and imbalanced ImageNet.
@@
 Figure 8: Visualisations of the learned $\lambda$ and $\kappa$ parameters for balanced ImageNet1K (IN-IK) training in blue, and imbalanced ImageNet-LT (IN-LT) training in blue.
-0.C.6.1 Visualisations on Imagenet-LT
+C-D1 Visualisations on Imagenet-LT
 We further show more qualitative results on ImageNet-LT with ResNet50 backbone.
-On the left subfigure, we show the model’s highest prediction marked with F,C,R that stands for frequent, common and rare class respectively and the Grad-cam activation [75].
+On the left subfigure, we show the model’s highest prediction marked with F,C,R that stands for frequent, common and rare class respectively and the Grad-cam activation [selvaraju2017grad].
 On the right subfigure, we show the last layer’s channel attention signal and its corresponding entropy denoted with (E).
@@
 Our method produces attention signals that have significantly larger entropy than the baseline for both frequent and rare classes.
-0.C.7 Calibration results
+Figure 10: Calibration results using ResNets on ImageNet-LT.
+SE (left) is underconfident, i.e., its confidence scores are lower than its actually accuracy due to over-regularisation.
+Our APA* (right) reduces the ECE and makes more accurate predictions with higher confidence than SE.
+C-E Calibration results
 Calibration is an important property of models, since it reassures that the confidence of the prediction matches the actual accuracy.
@@
 When APA* is applied, it reduces the Expected Calibration Error (ECE) as shown in Figure 10 (right-subfigure) for all backbones.
-Figure 8: Calibration results using ResNets on ImageNet-LT.
-SE (left) is underconfident, i.e., its confidence scores are lower than its actually accuracy due to over-regularisation.
-Our APA* (right) reduces the ECE and makes more accurate predictions with higher confidence than SE.
-0.C.8 Next textual token prediction experiment
-We perform one preliminary next-token prediction experiment using GPT2 [70] and the FineWeb-Edu [59] subset that contains 10 billion GPT2 tokens.
-The model is based on the GPT2 smallest architecture, which contains 117M parameters, and the code implementation follows [45].
-We train the GPT2 model for one epoch, using 8 V100 GPUs, a total batch size of 0.5M tokens, learning rate $46e-$ 4 and Adam optimizer [48] with momentum.
-We test the model on the HellaSwag benchmark [105] using zero-shot evaluation.
-To apply AGLU with GPT2, we simply switch the GELU activation with AGLU inside all MLP layers of the transformer.
-As the results suggest in Table 15, our AGLU increases the performance of GPT2 by $0.4\%$, showing that AGLU could be a good alternative for text-classification.
-Table 15: Comparative results using GPT2 smallest model and HellaSwag benchmark.
-Method Accuracy
-GELU 31.0
-AGLU 31.4